Jefferson Lab HALL A Logbook for March 12 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 368794 03/31/12 23:54 Yuxiang zhao [target]summary for the swing shift
  • 368793 03/31/12 23:53 Chao Scaler plots for 3rd arm run 40600 with slow raster size 1.8cm
  • 368792 03/31/12 23:46 Chao Sport++ for run 3556 with Slow Raster size 1.8cm (2)
  • 368791 03/31/12 23:45 Chao Sport++ for run 3556 with Slow Raster size 1.8cm (1)
  • 368775 03/31/12 23:23 Chao Asked MCC to send 50nA CW beam
  • 368768 03/31/12 22:47 johans Deadtime high, restart
  • 368767 03/31/12 22:45 johans Deadtime too high, restarted
  • 368766 03/31/12 22:44 Yuxiang zhao [target]dilution run summary
  • 368759 03/31/12 22:29 Chao Third arm rate at dummy target
  • 368748 03/31/12 22:11 Chao Third arm rate on carbon target
  • 368742 03/31/12 21:58 Chao Thrid arm rate at empty target
  • 368738 03/31/12 21:43 Chao Charge asymmetry plots for run 3549
  • 368737 03/31/12 21:42 Chao Charge asymmetry for run 3549
  • 368736 03/31/12 21:27 Chao Move HRS setting to 1.441
  • 368730 03/31/12 20:59 happex_auto happex bcm && bpm && lumi
  • 368729 03/31/12 20:54 Chao Spot++ for run 3548 (2)
  • 368728 03/31/12 20:54 Chao spot++ for run 3548 (1)
  • 368727 03/31/12 20:50 Chao Bleed through test
  • 368723 03/31/12 20:30 M. Yurov [target] Polarized Top Cup
  • 368721 03/31/12 20:26 D. Keller [Target] Polarization of Top Target
  • 368720 03/31/12 20:07 mhuang db_run.dat info lost
  • 368719 03/31/12 20:00 johans run numbers
  • 368718 03/31/12 19:48 johans BPM and Raster
  • 368717 03/31/12 19:47 D. Keller [Target] New Baseline
  • 368716 03/31/12 19:43 johans Raster/BPM L
  • 368715 03/31/12 19:43 A. Camsonne Could you specify run number ?
  • 368714 03/31/12 18:58 jmaxwell [Target] Anneal Done
  • 368713 03/31/12 18:06 adaq Replay error
  • 368710 03/31/12 18:00 jmaxwell [Target] Starting Anneal
  • 368707 03/31/12 17:33 Kalyan Thirdarm rates with 70nA
  • 368705 03/31/12 17:29 johans 3546Coda crash
  • 368698 03/31/12 16:51 johans No end of run 3544
  • 368682 03/31/12 16:11 T. Badman Day shift summary
  • 368679 03/31/12 16:05 jmaxwell Slow Raster ADC Auto Post Request
  • 368676 03/31/12 16:01 R. Michaels duplicate coda files
  • 368665 03/31/12 15:41 adaq IHWP status is IN
  • 368663 03/31/12 15:41 T. Badman HWP changed to IN
  • 368661 03/31/12 15:40 pzhu spot++ and spoth++ script update
  • 368655 03/31/12 14:55 pzhu spot++ for the newest run
  • 368654 03/31/12 14:48 T. Badman spot++
  • 368648 03/31/12 14:23 T. Badman late- MCC finished tuning beam
  • 368645 03/31/12 14:11 T. Badman Spot++ after tune
  • 368637 03/31/12 13:57 jmaxwell [Target] Moved to Bottom
  • 368636 03/31/12 12:38 T. Badman harp scan
  • 368635 03/31/12 12:29 jie Target moved from top to home for harp scan
  • 368631 03/31/12 12:10 T. Badman scaler rates
  • 368618 03/31/12 11:36 jmaxwell [Target] Take Advantage of Unstable Beam
  • 368617 03/31/12 11:32 T. Badman slow raster ADC plot
  • 368616 03/31/12 11:22 T. Badman Beam position moved
  • 368603 03/31/12 10:39 T. Badman Unstable beam most of the day
  • 368594 03/31/12 10:14 T. Badman HWP moved to OUT
  • 368592 03/31/12 10:13 adaq IHWP status is OUT
  • 368584 03/31/12 09:36 jie target movement for dilution run
  • 368583 03/31/12 09:33 T. Badman Dilution run summary
  • 368570 03/31/12 09:14 T. Badman Runs 22561/3528/40566 are C12 dilution runs
  • 368556 03/31/12 08:16 T. Badman Hall A Tools
  • 368555 03/31/12 08:12 pzhu owl shift summary
  • 368550 03/31/12 07:35 pzhu run 22558,225559,22555,22554,22553,22552,22551 is production run
  • 368546 03/31/12 06:52 R. Michaels problem fix : no R-HRS ev140 scalers in data
  • 368539 03/31/12 06:06 adaq IHWP status is IN
  • 368538 03/31/12 06:06 adaq IHWP status is IN
  • 368534 03/31/12 05:59 happex_auto happex bcm && bpm && lumi
  • 368533 03/31/12 05:47 pzhu bleed-through measurement
  • 368520 03/31/12 03:39 pzhu 22551 and 22552 is production run
  • 368510 03/31/12 03:09 pzhu spot++ check for 1.5V and 1.1V
  • 368509 03/31/12 03:09 Kalyan Re: thirdarm rate
  • 368502 03/31/12 02:44 Kalyan Third arm rate increase after Moller
  • 368501 03/31/12 02:40 jmaxwell [Target] OVC Pressure Access
  • 368500 03/31/12 02:33 A. Camsonne Gas for MRPC moved in the Hall
  • 368499 03/31/12 02:24 R. Michaels [target] recovery of target, status
  • 368498 03/31/12 02:20 jmaxwell [Target] Magnet Up
  • 368497 03/31/12 02:05 adaq IHWP status is OUT
  • 368495 03/31/12 01:54 jmaxwell Septa and Target Magnet Off, Locked
  • 368494 03/31/12 01:45 jmaxwell [Target] EIP Reset
  • 368493 03/31/12 01:34 R. Michaels controlled access to check target, P0 change
  • 368492 03/31/12 01:27 Chao Septa is off
  • 368487 03/31/12 01:08 R. Michaels [target] top tgt put in, uwaves on
  • 368482 03/31/12 00:59 R. Michaels various target crises
  • 368481 03/31/12 00:58 jmaxwell [Target] EIP Read Problems
  • 368476 03/31/12 00:54 R. Michaels [target] at dummy target, then home
  • 368467 03/31/12 00:32 R. Michaels [target] frequency 140 ?
  • 368466 03/31/12 00:31 R. Michaels [target] move to C12
  • 368462 03/31/12 00:08 Ole Hansen Swing Shift Summary
  • Friday

  • 368456 03/30/12 23:31 Ole Hansen spot++ screen 3
  • 368455 03/30/12 23:31 Ole Hansen spot++ during production
  • 368443 03/30/12 22:33 Jixie, Min, beam position dependence
  • 368437 03/30/12 22:03 Ole Hansen Hall A Tools
  • 368435 03/30/12 22:03 Ole Hansen Target polarization
  • 368433 03/30/12 21:57 Ole Hansen Start production
  • 368432 03/30/12 21:55 Ole Hansen Target in position, polarizing ...
  • 368431 03/30/12 21:52 Ole Hansen IHWP IN
  • 368430 03/30/12 21:49 Ole Hansen Target looks well centered
  • 368429 03/30/12 21:48 Ole Hansen spot_L run 3507
  • 368427 03/30/12 21:46 adaq IHWP status is IN
  • 368422 03/30/12 21:46 Ole Hansen spot_R for run 22358
  • 368420 03/30/12 21:43 Jixie Spot_L 3506
  • 368417 03/30/12 21:41 Ole Hansen Target centering
  • 368414 03/30/12 20:35 Ole Hansen No usable beam yet
  • 368413 03/30/12 20:06 Ole Hansen Target TE done
  • 368411 03/30/12 20:02 T. Badman TE results
  • 368410 03/30/12 20:00 Chao Third arm HV update
  • 368407 03/30/12 19:15 jmaxwell [Target] Access for RF Error
  • 368405 03/30/12 18:55 Ole Hansen Controlled access
  • 368403 03/30/12 18:34 Ole Hansen Target problems
  • 368402 03/30/12 18:28 Sasha Moller measurement results
  • 368401 03/30/12 18:20 R. Zielinski ROC 3 Error Message
  • 368399 03/30/12 18:05 Ole Hansen Moller measurement done
  • 368398 03/30/12 17:50 happex_auto happex bcm && bpm && lumi
  • 368397 03/30/12 16:49 Sasha IHWP is OUT
  • 368396 03/30/12 16:48 adaq IHWP status is OUT
  • 368394 03/30/12 16:15 A. Camsonne HAPPEX DAQ and HARP run 3376
  • 368393 03/30/12 16:01 kamyers Day shift summary
  • 368392 03/30/12 15:52 A. Camsonne Spot_L 3376
  • 368391 03/30/12 15:44 jmaxwell [Target] eLog Discrepancy
  • 368390 03/30/12 15:41 Solvignon Target screenshot
  • 368389 03/30/12 15:14 jmaxwell [Target] eLog Discrepancy
  • 368388 03/30/12 13:58 kamyers Bleed-through measurement - 0.5%
  • 368387 03/30/12 13:55 kamyers LHRS and septum back to 1.792 GeV setting
  • 368386 03/30/12 13:51 solvignon Move target home for Moller measurement
  • 368383 03/30/12 13:44 solvignon Move target to bottom cell
  • 368380 03/30/12 13:27 jmaxwell [Target] Anneal Done
  • 368379 03/30/12 13:27 Chao g2p softwares on adaq1
  • 368378 03/30/12 13:21 kamyers LHRS to elastic 2.228 GeV setting
  • 368377 03/30/12 12:01 kamyers Beam off for target anneal
  • 368360 03/30/12 11:05 kamyers IHWP IN, passive HWP OUT
  • 368353 03/30/12 10:50 kamyers IHWP OUT, but passive IHWP changed to IN
  • 368348 03/30/12 10:37 kamyers Back to 50 nA (about 15 minutes ago)
  • 368345 03/30/12 10:35 kamyers Beam positions slightly off nominal since 18:00 last night
  • 368341 03/30/12 10:03 kamyers Beam current 50 nA -> 65 nA
  • 368340 03/30/12 09:58 yawei DAQs snapshot
  • 368327 03/30/12 09:30 Solvignon Screenshot: Target control
  • 368326 03/30/12 09:15 jmaxwell [Target] Main Flow Alarm Change
  • 368322 03/30/12 08:59 kamyers FR to 2x2, Slow raster on
  • 368321 03/30/12 08:59 kamyers 1.792 setting
  • 368320 03/30/12 08:20 kamyers Septa for 1.792 setting
  • 368319 03/30/12 08:16 R. Zielinski Owl Shift Summary
  • 368314 03/30/12 08:02 M. Yurov [target] Moved target
  • 368309 03/30/12 07:46 M. Yurov [target] Moved target
  • 368308 03/30/12 07:45 M. Yurov [target] Typo in the log entry 368304
  • 368305 03/30/12 07:40 M. Yurov [target] Moved target
  • 368304 03/30/12 07:38 M. Yurov [target] Main flow alarm. Missing events in PDP
  • 368301 03/30/12 07:24 R. Zielinski Finished 1.927 GeV Production. Switching to dilution
  • 368297 03/30/12 07:06 M. Yurov [target] PDP is missing events
  • 368282 03/30/12 05:41 happex_auto happex bcm && bpm && lumi
  • 368274 03/30/12 04:59 R. Zielinski Hall C Changing HWP to OUT
  • 368264 03/30/12 03:43 R. Zielinski Full Replays of all 2.072 GeV data
  • 368233 03/30/12 01:04 R. Zielinski Hall C changing Half-Wave Plate to IN
  • 368232 03/30/12 01:02 R. Zielinski Killed Third Arm CODA
  • 368226 03/30/12 00:51 adaq IHWP status is IN
  • 368216 03/30/12 00:21 R. Zielinski Hall A Tools
  • 368215 03/30/12 00:02 T. Badman swing shift summary
  • Thursday

  • 368208 03/29/12 23:46 Aniol target - top cup calibration constant
  • 368207 03/29/12 23:39 Aniol target - tgt elog discrepancy
  • 368206 03/29/12 23:21 Aniol target - top cup baseline
  • 368204 03/29/12 23:11 Aniol target - move to top cup for TE
  • 368194 03/29/12 22:40 siqbal R-arm run #22511, change in comments
  • 368188 03/29/12 21:48 siqbal run #s 3475 & 40525 junk
  • 368181 03/29/12 21:28 Aniol target - bottom cell moved in
  • 368180 03/29/12 21:27 T. Badman changing HRS momentum setting
  • 368179 03/29/12 21:25 T. Badman Dilution Run Summary
  • 368160 03/29/12 20:42 siqbal L-arm run #3470, change in comments
  • 368153 03/29/12 20:26 Aniol target-moved to HOME tgt
  • 368146 03/29/12 20:08 Aniol target - moved to cross hair dummy
  • 368137 03/29/12 19:46 Aniol target - moved to c12_40
  • 368127 03/29/12 18:47 Chao End of run information for run 22501
  • 368121 03/29/12 18:35 Jixie Right dipole ramping speed
  • 368120 03/29/12 18:35 Chao Third arm HV update
  • 368116 03/29/12 18:00 Aniol target - TE calibration
  • 368115 03/29/12 17:31 slifer Automated Run Summaries
  • 368114 03/29/12 17:24 aniol target moved to HOME
  • 368113 03/29/12 17:23 Jixie Zhang Tools to calculate septum current
  • 368111 03/29/12 16:55 T. Badman Hall A General Tools
  • 368108 03/29/12 16:46 siqbal running TA cosmics
  • 368107 03/29/12 16:43 T. Badman beam down for 1 hour
  • 368105 03/29/12 16:09 T. Badman HWP changed
  • 368104 03/29/12 16:05 adaq IHWP status is OUT
  • 368102 03/29/12 15:59 yawei Day Shift Summary
  • 368101 03/29/12 15:41 johans Beamtrip
  • 368097 03/29/12 15:36 pzhu left arm upstream bcm scaler abnormal
  • 368090 03/29/12 14:58 Kalyan Third arm high voltages
  • 368083 03/29/12 14:41 brads Rotatable HWP changes during day shift?
  • 368082 03/29/12 14:22 jmaxwell [Target] New Baseline on Bottom
  • 368071 03/29/12 13:26 johans No start entry 40509
  • 368055 03/29/12 11:52 pzhu half-wave plate variable not in raw data
  • 368048 03/29/12 11:31 Chao Charge Asymmetry check
  • 368041 03/29/12 10:47 yawei Hall A tools
  • 368032 03/29/12 10:14 adaq Incorrect Run information and scalers
  • 368031 03/29/12 10:04 yawei run# 3455 no start of run
  • 368024 03/29/12 09:50 happex_auto happex bcm && bpm && lumi
  • 368014 03/29/12 09:30 yawei MCC starting to tune beam into Hall A
  • 368013 03/29/12 08:00 M. Cummings Owl Shift Summary
  • 368012 03/29/12 07:03 M. Yurov [target] Moved target back to NH3 Bottom Cup
  • 368011 03/29/12 06:39 M. Cummings Changed septum settings
  • 368009 03/29/12 06:37 M. Cummings Changing LHRS and RHRS dipole to 2.072
  • 368000 03/29/12 06:28 M. Yurov [target] Moved target
  • 367995 03/29/12 06:19 M. Yurov [target] Moved target
  • 367970 03/29/12 05:55 M.Yurov [target] Moved target
  • 367966 03/29/12 05:50 happex_auto happex bcm && bpm && lumi
  • 367964 03/29/12 05:49 M. Yurov [target] Moved target
  • 367957 03/29/12 05:40 M. Yurov [target] Moved target
  • 367956 03/29/12 05:32 M. Cummings Done with elastic setting - starting dilution runs
  • 367953 03/29/12 05:28 happex_auto happex bcm && bpm && lumi
  • 367952 03/29/12 05:25 happex_auto happex bcm && bpm && lumi
  • 367950 03/29/12 05:06 M. Cummings T2 rate high
  • 367940 03/29/12 04:17 M. Cummings Septum Strip Charts
  • 367939 03/29/12 04:12 pzhu changed ped for roc6,slot 12
  • 367938 03/29/12 04:06 M. Yurov [target] Nose level PID loop
  • 367936 03/29/12 04:03 M. Cummings RHRS rate back to normal
  • 367931 03/29/12 03:53 pzhu one sieve slit is in before
  • 367930 03/29/12 03:51 M. Cummings Back in beam permit, asked for 50 nA
  • 367929 03/29/12 03:46 pzhu sieve slit is half in before
  • 367928 03/29/12 03:44 M. Yurov [target]
  • 367927 03/29/12 03:37 M. Yurov [target] Change target position.
  • 367926 03/29/12 03:35 M. Yurov [target] Change target position.
  • 367924 03/29/12 03:33 A. Camsonne HRS angles
  • 367923 03/29/12 03:27 M. Cummings Called for escorted access to check sieve slit
  • 367922 03/29/12 03:26 pzhu center beam position
  • 367917 03/29/12 03:20 M. Cummings Hall C changing HWP to OUT
  • 367913 03/29/12 03:01 M. Cummings RHRS deadtime - other RHRS component windows
  • 367912 03/29/12 02:59 M. Cummings high deadtime on RHRS
  • 367906 03/29/12 02:30 A. Camsonne Right HRS spot_R not working
  • 367903 03/29/12 02:18 A. Camsonne BPM 1H05B saturated for previous runs
  • 367898 03/29/12 02:10 A. Camsonne BPM set to dual gain
  • 367894 03/29/12 01:55 Chao dp plots check for HRS
  • 367893 03/29/12 01:54 M. Cummings Called for 50nA - still working on low RHRS rates
  • 367892 03/29/12 01:39 M. Cummings RHRS rate low
  • 367886 03/29/12 01:35 A. Camsonne BPM receivers settings auto-gain
  • 367882 03/29/12 01:33 M. Yurov [target] Polarization
  • 367881 03/29/12 00:55 A. Camsonne Hall A tungsten calorimeter run 75 nA
  • 367873 03/29/12 00:20 mhuang Sieve is OUT!!!
  • 367872 03/29/12 00:20 A. Camsonne BCM calibration tungsten calorimeter run 75 nA
  • 367870 03/29/12 00:17 M. Cummings Hall A Tools
  • 367869 03/29/12 00:15 mhuang Sieve moved in
  • 367868 03/29/12 00:14 M. Yurov [target] New Baseline
  • 367867 03/29/12 00:12 M. Cummings Going back to beam permit
  • Wednesday

  • 367866 03/28/12 23:55 Jixie, Min, optics summary of 2.5T, 90deg target field and 2.254 GeV beam, 40-32-16 right septum
  • 367865 03/28/12 23:39 R. Zielinski Swing Shift Summary
  • 367864 03/28/12 23:35 Jixie Escort acess to move sieve out
  • 367861 03/28/12 23:35 Jixie Finished optics
  • 367860 03/28/12 23:32 siqbal spot_L 3436 & spot_R 22469
  • 367857 03/28/12 23:25 Jixie Slow raster on, 1.4V
  • 367852 03/28/12 23:13 Jixie Turn on faster raster 2x2
  • 367850 03/28/12 23:11 Jixie Finish delta scan. Set both arms to H elastic (2.228)
  • 367848 03/28/12 23:05 siqbal liquid helium
  • 367843 03/28/12 22:58 Jixie left delta to -1%
  • 367841 03/28/12 22:56 Jixie Should be -3%, not 3%
  • 367839 03/28/12 22:52 Production runplan for tonight
  • 367837 03/28/12 22:48 A. Camsonne Spot L Run 3430
  • 367834 03/28/12 22:42 Jixie left delta =-2%
  • 367827 03/28/12 22:27 Jixie left delta = 3%
  • 367821 03/28/12 22:09 adaq LEFT -4% right 1%
  • 367816 03/28/12 21:50 Jixie Cycle Left Q2,Q3, set to delta=4%
  • 367812 03/28/12 21:35 Chao End of Run information of run 3423
  • 367800 03/28/12 20:58 mhuang lhrs 3% snapshots
  • 367793 03/28/12 20:40 mhuang Lhrs 2%, rhrs 2% snapshots
  • 367790 03/28/12 20:32 mhuang Start delta scan, left 1%, right 2%
  • 367781 03/28/12 20:13 siqbal L arm coda crash
  • 367774 03/28/12 19:57 siqbal L arm coda crash
  • 367771 03/28/12 19:49 Jixie Move beam back to nominal (0,0)
  • 367766 03/28/12 19:34 Jixie change beam position to (-4,0)
  • 367761 03/28/12 19:21 Jixie Change beam position to (-4,-4)
  • 367758 03/28/12 19:17 siqbal R-arm coda crash
  • 367754 03/28/12 19:07 Jixie Change beam position to (0,-4)
  • 367747 03/28/12 18:51 Jixie Change beam position to (0,4)
  • 367746 03/28/12 18:46 Jixie Hall C change HWP
  • 367733 03/28/12 18:31 A. Camsonne BCM receiver slope updated
  • 367726 03/28/12 18:16 Jixie Ramp septum to 725.4(left) and 806.6(right)
  • 367725 03/28/12 18:14 A. Camsonnne MRPC work
  • 367724 03/28/12 18:14 Jixie beam is back, continue optics
  • 367723 03/28/12 17:57 pzhu bpm receiver setting backup
  • 367722 03/28/12 17:51 Jack(Techs) Septa
  • 367717 03/28/12 16:56 R. Zielinski Escorted access for Jack
  • 367716 03/28/12 16:31 Jixie Ramp down septa to zero
  • 367715 03/28/12 16:26 R. Zielinski Hall A Tools
  • 367714 03/28/12 16:23 jmaxwell [Target] Ramping Down Magnet
  • 367705 03/28/12 16:03 Ole Hansen Day Shift Summary
  • 367704 03/28/12 16:02 Jixie Change right septum to 810A
  • 367695 03/28/12 15:46 Jixie Change right septum to 802.6
  • 367690 03/28/12 15:34 Jixie change right septum to 790.5A
  • 367681 03/28/12 15:20 Sasha Rates for runs 3400 and 22437
  • 367680 03/28/12 15:19 JIXIE change right septum to 798.5
  • 367675 03/28/12 15:11 mhuang Right septum current 806.0->806.6
  • 367672 03/28/12 15:10 Jack(Techs) Correct channel for Left Septum power supply voltage output
  • 367671 03/28/12 15:09 jmaxwell [Target] Low Nose Level Alarms Off for Low Helium Running
  • 367652 03/28/12 14:42 Kalyan Re: EPICS channels for the Septa
  • 367651 03/28/12 14:23 Ole Hansen No beam
  • 367646 03/28/12 14:21 Ole Hansen Commencing optics studies
  • 367645 03/28/12 14:20 Jixie Optics start, tools window snapshot
  • 367644 03/28/12 14:20 Ole Hansen Left septum reset
  • 367643 03/28/12 14:16 Sasha IHWP is OUT
  • 367642 03/28/12 14:15 Jixie Left septum is back
  • 367641 03/28/12 14:10 pzhu slow raster size calibration
  • 367640 03/28/12 14:09 Ole Hansen Still investigating septum issues (if any)
  • 367635 03/28/12 14:03 brads glitch in MBIGBOX variable?
  • 367634 03/28/12 14:00 Ole Hansen Installed ROOT & analyzer on adaq1
  • 367630 03/28/12 13:50 Jixie septum tripped
  • 367629 03/28/12 13:44 Ole Hansen Left Septum trip
  • 367628 03/28/12 13:44 pzhu center beam position
  • 367626 03/28/12 13:38 adaq center beam position check
  • 367625 03/28/12 13:37 Ole Hansen Asking for beam back, none yet
  • 367624 03/28/12 13:36 Ole Hansen Preparing for optics
  • 367623 03/28/12 13:36 Ole Hansen Target centering finished, starting optics
  • 367622 03/28/12 13:32 jmaxwell [Target] Move to Carbon 40 mil, Closed Run Valve
  • 367607 03/28/12 13:16 Ole Hansen Beam is back
  • 367606 03/28/12 12:58 pzhu fast raster calibration
  • 367605 03/28/12 12:56 Ole Hansen No beam, possibly for long while
  • 367584 03/28/12 11:44 pzhu harp scan
  • 367541 03/28/12 09:09 Ole Hansen Tune beam
  • 367540 03/28/12 08:46 Ole Hansen No beam for ca 30 minutes
  • 367539 03/28/12 08:45 Ole Hansen Hall C change IHWP to IN around 8:35
  • 367538 03/28/12 08:43 Ole Hansen Beam is back
  • 367537 03/28/12 08:34 pzhu no beam in one hour
  • 367536 03/28/12 08:28 mhuang Right septum current scan with harp scan
  • 367535 03/28/12 08:25 Ole Hansen No beam
  • 367534 03/28/12 08:24 Ole Hansen Day shift plan
  • 367529 03/28/12 08:10 D. Keller [Target] Moved C12-40
  • 367528 03/28/12 08:06 mhuang Owl shift summary
  • 367523 03/28/12 07:54 pzhu harp scan
  • 367522 03/28/12 07:53 pzhu beam spot in 50nA
  • 367509 03/28/12 07:27 pzhu bpm in high current
  • 367504 03/28/12 07:22 pzhu raster shape in bpm
  • 367500 03/28/12 07:19 pzhu raster shape in bpm
  • 367492 03/28/12 07:12 pzhu raster shape in BPM
  • 367489 03/28/12 07:09 pzhu raster shape in bpm
  • 367482 03/28/12 07:03 pzhu raster shape in BPM
  • 367473 03/28/12 06:56 pzhu raster shape in bpm
  • 367470 03/28/12 06:53 pzhu raster shape in BPM
  • 367465 03/28/12 06:49 pzhu raster shape in BPM
  • 367458 03/28/12 06:43 pzhu raster shape in BPM
  • 367453 03/28/12 06:37 pzhu slow raster shape in bpm
  • 367448 03/28/12 06:30 pzhu Re:mismatched for hac_bcm_average and IBC1H00MAGA
  • 367446 03/28/12 06:27 pzhu raster shape in bpm for low current pulser beam
  • 367438 03/28/12 06:14 pzhu mismatched for hac_bcm_average and IBC1H00MAGA
  • 367409 03/28/12 04:38 D. Keller Hall C halfwave plate
  • 367405 03/28/12 04:01 mhuang Escorted access into Hall
  • 367391 03/28/12 02:34 Jechiel Scalers
  • 367390 03/28/12 02:33 pzhu reset TS0 crate
  • 367389 03/28/12 02:17 Jechiel R-HRS scalers
  • 367388 03/28/12 02:11 mhuang Hall C change IHWP to be OUT around 0:06
  • 367384 03/28/12 01:45 mhuang MCC still trying to tune beam into FZ magnet
  • 367382 03/28/12 00:08 M. Cummings Swing Shift Summary
  • 367381 03/28/12 00:00 Aniol target - buffer level fluctuates
  • Tuesday

  • 367380 03/27/12 23:33 M. Cummings Left and Right septum strip charts
  • 367379 03/27/12 23:10 M. Cummings Hall to controlled access so MCC can check out power supply for FZ magnet
  • 367378 03/27/12 23:05 A. Camsonne Added raw ADCs from BPM 1H05A and 1H05B receiver in fastEpicsLogger
  • 367377 03/27/12 23:00 Kalyan Re: EPICS channels for the Septa
  • 367376 03/27/12 22:54 M. Cummings MCC still having FSD problems, fz magnet is down
  • 367375 03/27/12 21:58 slifer runplan update
  • 367374 03/27/12 21:52 M. Cummings Beam tuning troubles
  • 367373 03/27/12 21:20 M. Cummings MCC tuning beam into hall
  • 367372 03/27/12 21:15 Jack(Techs) EPICS channels for the Septa
  • 367371 03/27/12 21:13 M. Cummings Septum ramped up
  • 367370 03/27/12 21:02 J Mulholland Level probe may stick
  • 367369 03/27/12 20:50 A. Camsonne MRPC DAQ installed
  • 367368 03/27/12 20:49 Aniol target - buffer level?
  • 367367 03/27/12 20:45 Jack(Techs) Revised Septa current limit
  • 367366 03/27/12 20:39 Kalyan Right septum limit and tonights optics plan
  • 367365 03/27/12 20:37 Aniol target - buffer dewar levels
  • 367364 03/27/12 20:23 M. Cummings Ramping up Left and Right dipoles to 2.251
  • 367363 03/27/12 20:00 M. Cummings Called for Sweep
  • 367362 03/27/12 19:56 mhuang Sieve is IN now
  • 367361 03/27/12 19:45 Jack(Techs) Septum magnets
  • 367360 03/27/12 18:34 Kalyan Problem running thirdarm daq on adaql2
  • 367354 03/27/12 15:54 Kalyan Status Update
  • 367349 03/27/12 15:09 Fansler Checklist
  • 367329 03/27/12 11:46 Kalyan Possibility to lock-up during swing shift - Please check you subsytems
  • Monday

  • 367305 03/26/12 14:41 slifer Runplan for 2.2 GeV
  • 367303 03/26/12 09:53 Kalyan Shifts Update: shifts canceled through Tue OWL, Tue DAY is on stand-by
  • 367296 03/26/12 09:14 pzhu happex ADC gain changed
  • 367295 03/26/12 09:04 R. Michaels to-do list for Hall A to come back
  • Sunday

    
    

    Saturday

  • 367293 03/24/12 23:55 pzhu second happex timing board installed in both arm's happex
  • Friday

  • 367291 03/23/12 15:59 brads IHWP status change alarm hander available
  • 367290 03/23/12 14:07 Chao Add one DISC to 3rd arm
  • 367286 03/23/12 12:09 A. Camsonne Those are BCM cables
  • Thursday

  • 367285 03/22/12 19:46 Kalyan g2p/gep status update and shift schedule
  • 367282 03/22/12 18:41 R. Zielinski CODA Configurations bypassing @bobfile
  • 367268 03/22/12 17:58 pzhu calibrated bpm
  • 367262 03/22/12 12:44 brads Cable run off the split to John's box
  • 367261 03/22/12 12:31 A. Camsonne Those are BCM cables
  • 367259 03/22/12 12:00 brads Uneeded BPM split upstream of Musson box?
  • 367257 03/22/12 09:39 Kalyan g2p/gep Shift Update: Thurs. SWING and Fri. OWL are canceled, Fri DAY is on standby
  • 367255 03/22/12 00:06 pzhu changed the ADC pedestal for BPM ADC
  • Wednesday

  • 367254 03/21/12 23:41 pzhu connected bpm to 400ns ADC in left arm
  • 367239 03/21/12 19:10 Jack(Techs) BigBox "One Transistor Fault"
  • 367238 03/21/12 17:00 R. Michaels re: duplicate CODA files
  • 367237 03/21/12 14:12 M. Cummings Timing of T2
  • 367236 03/21/12 11:41 Jixie Zhang Survey results are available in g2p wiki
  • 367233 03/21/12 10:44 brads g2p/gep shift update: Shifts are canceled through Thursday DAY, Thursday SWING+ on standby
  • 367232 03/21/12 01:35 mhuang Preliminary optics dp calibration result of LHRS
  • Tuesday

  • 367231 03/20/12 19:06 R. Zielinski Added 1881M added to LHRS in ROC 4
  • 367230 03/20/12 19:05 R. Zielinski Added 1881M added to LHRS in ROC 4
  • 367221 03/20/12 18:35 R. Zielinski 1881 ADC added to ROC 6 in RHRS
  • 367217 03/20/12 18:06 Chao Left arm fast clock problem is fixed
  • 367210 03/20/12 16:15 pzhu bpm noise for different gain
  • 367192 03/20/12 14:53 brads g2p/gep Tuesday SWING, Wed OWL, Wed DAY cancelled; Wed SWING+ on standby (Brief status update)
  • 367191 03/20/12 14:30 T. Badman left arm fastclock count in ring
  • 367177 03/20/12 10:34 Chao Adjust 3rd arm HV with cosmic ray test
  • 367176 03/20/12 10:33 Kalyan Connected dry air flow to thirdarm dE plane PMTs
  • 367175 03/20/12 08:54 Andrew(techs LCW Water
  • Monday

  • 367173 03/19/12 14:01 brads gep/g2p update -- shifts canceled through Tuesday OWL; Tuesday DAY on standby
  • 367170 03/19/12 08:44 brads Hall Access -- keep it to a minimum please
  • Sunday

  • 367167 03/18/12 20:39 slifer run summary
  • 367166 03/18/12 18:11 Jack(Techs) Right Q3 current limit is now 1550 amps
  • 367162 03/18/12 07:09 pzhu Re:more duplicate CODA files
  • Saturday

  • 367159 03/17/12 22:49 Jack VDCs high voltage set to zero
  • 367158 03/17/12 22:20 Jack(Techs) More puzzle pieces
  • 367157 03/17/12 21:34 slifer elastic g2p production runs
  • 367156 03/17/12 21:33 slifer elastic g2p production runs
  • 367154 03/17/12 16:17 Kalyan Shifts cancelled for this weekend
  • 367153 03/17/12 15:28 Jack (Techs) Magnets ramped down
  • 367142 03/17/12 14:47 Yuncheng Stop Blowers and Open Valves
  • 367141 03/17/12 14:25 R. Michaels more duplicate CODA files
  • 367140 03/17/12 13:44 yawei Day shift summary
  • 367138 03/17/12 13:21 Folts VESDA/tech
  • 367137 03/17/12 13:09 Higinbotham All The Pieces of the Puzzle
  • 367136 03/17/12 12:32 Jack(Techs) Right Septum
  • 367123 03/17/12 11:32 jmaxwell [Target] Level Alarm
  • 367122 03/17/12 11:19 Kalyan Update
  • 367121 03/17/12 11:03 yawei Hall C changing half wave plate from out->in
  • 367120 03/17/12 10:57 yawei Hall is in restricted access
  • 367119 03/17/12 10:51 Yuncheng Target magnet ramping down
  • 367118 03/17/12 10:47 jmaxwell [Target] Magnet Ramped Down, Leads Locked
  • 367116 03/17/12 10:32 mhuang No gap
  • 367115 03/17/12 10:06 yawei ramp down the target magnet
  • 367114 03/17/12 10:03 Yuncheng Target Alarm, Nose level, Magnet ramping down
  • 367113 03/17/12 09:38 Kalyan Both septa magnets are down
  • 367112 03/17/12 09:26 yuxiang zhao junk run for L:3295 R:22375
  • 367111 03/17/12 09:25 mhuang Gap in the middle
  • 367110 03/17/12 09:25 yawei fire alarm and septa trip
  • 367108 03/17/12 09:16 yawei right septa tripped
  • 367095 03/17/12 08:45 yuxiang zhao screen shot after setting prescale 4 on LHRS & prescale 2 on RHRS
  • 367093 03/17/12 08:40 yuxiang zhao prescale setting
  • 367089 03/17/12 08:18 yuxiang zhao LHRS run number 3292 pr1 RHRS run number 22372 preshower1
  • 367088 03/17/12 08:06 mhuang Owl shift summary
  • 367076 03/17/12 07:16 hkkang s shot 22371
  • 367075 03/17/12 07:14 mhuang Hall C changing IHWP to be OUT
  • 367074 03/17/12 07:13 hkkang s shot 3291
  • 367071 03/17/12 06:58 mhuang beam charge asymmetry check-TA
  • 367068 03/17/12 06:52 mhuang beam charge asymmetry check
  • 367067 03/17/12 06:49 hkkang snap for 22370RHRS
  • 367064 03/17/12 06:41 hkkang snap of run3290
  • 367054 03/17/12 06:13 D. Keller [Target] polarization
  • 367046 03/17/12 05:26 hkkang snap shot (run3285,22366)
  • 367045 03/17/12 05:19 mhuang Bleed through result
  • 367042 03/17/12 05:09 mhuang Bleed through measurement-II
  • 367037 03/17/12 04:48 hkkang snan shot 3284-22365
  • 367030 03/17/12 04:30 hkkang snap shot for production run
  • 367029 03/17/12 04:15 pzhu lumi detector after changing ADC gain
  • 367020 03/17/12 03:50 pzhu scaler rate
  • 367019 03/17/12 03:48 mhuang Bleed through measurement
  • 367018 03/17/12 03:45 pzhu slow raster check after first production
  • 366998 03/17/12 03:11 mhuang Hall C changing IHWP to be IN
  • 366997 03/17/12 03:02 adaq raster size calibration
  • 366994 03/17/12 02:59 pzhu center beam position double check
  • 366993 03/17/12 02:56 mhuang Hall A Tools
  • 366992 03/17/12 02:54 A. Camsonne EPICS position change with current
  • 366963 03/17/12 02:11 M. Cummings Run number for dilution runs
  • 366957 03/17/12 01:59 pzhu lumi hv is too small
  • 366951 03/17/12 01:49 M. Cummings Run numbers for Current Scan/Rate Study
  • 366915 03/17/12 00:12 M. Cummings scaler rates for 100nA
  • 366910 03/17/12 00:05 Ole Hansen Swing Shift Sumamry
  • Friday

  • 366904 03/16/12 23:54 M. Cummings scaler rates for 75 nA
  • 366903 03/16/12 23:52 A. Camsonne HARP scans from BPM calibration
  • 366894 03/16/12 23:33 M. Cummings scaler rates for 50 nA
  • 366885 03/16/12 23:16 M. Cummings scaler rates for 25nA
  • 366881 03/16/12 23:07 Ole Hansen Hall C changes IHWP
  • 366877 03/16/12 22:57 M. Cummings scaler rates for 15nA
  • 366874 03/16/12 22:47 R. Zielinski Beginning DAQ Current/Deadtime Scan
  • 366873 03/16/12 22:44 Ole Hansen Begin dilutions runs
  • 366870 03/16/12 22:38 Ole Hansen BPM & BCM calibrations done
  • 366869 03/16/12 22:38 kamyers [Target] Target moved to Carbon
  • 366868 03/16/12 22:25 A. Camsonne BCM calibration tungsten calorimeter run started
  • 366864 03/16/12 22:13 pzhu bpm information in different filter,different div
  • 366853 03/16/12 21:59 Kalyan HV change on Lumi 1 and 2
  • 366824 03/16/12 21:18 A. Camsonne BCM receiver setting 100 nA
  • 366820 03/16/12 21:10 A. Camsonne HARP scan new position
  • 366816 03/16/12 20:52 A. Camsonne HARP scan for position check
  • 366811 03/16/12 20:44 pzhu 25nA bpm
  • 366798 03/16/12 20:25 A. Camsonne Run 50 nA - Mag Filter 8 and div to 1
  • 366785 03/16/12 20:12 A. Camsonne Going to 100 nA CW
  • 366780 03/16/12 20:11 A. Camsonnne HARP scan
  • 366779 03/16/12 20:03 A. Camsonne Beam is back
  • 366778 03/16/12 19:15 M. Cummings RHRS acceptance
  • 366777 03/16/12 19:15 A. Camsonne Beam down
  • 366776 03/16/12 19:08 pzhu bpm information in different current
  • 366771 03/16/12 19:02 A. Camsonne HARP scan 1H05A(2,1.48)
  • 366770 03/16/12 18:55 A. Camsonne HARP scan
  • 366763 03/16/12 18:26 A. Camsonne Next run with with Mag Div 1 on BPM
  • 366756 03/16/12 18:21 A. Camsonne BPM receiver settings for 50 nA or more
  • 366751 03/16/12 18:12 A. Camsonne HARP scan initial position
  • 366741 03/16/12 17:47 kamyers [Target] target moved out of beam
  • 366740 03/16/12 17:43 Ole Hansen RHRS rate problem apparently fixed
  • 366735 03/16/12 17:34 Ole Hansen Beam positions adjusted
  • 366730 03/16/12 17:19 M. Cummings Signal from patch panel Betty C8 on RHRS
  • 366729 03/16/12 17:07 mhuang runnumber of the first entry
  • 366728 03/16/12 17:04 M. Cummings Sieve slits are OUT
  • 366727 03/16/12 16:51 mhuang acceptance plots of sieve IN run
  • 366726 03/16/12 16:40 Ole Hansen Controlled access
  • 366721 03/16/12 16:30 Ole Hansen Choppy beam, brief run taken
  • 366715 03/16/12 16:12 Ole Hansen Beam imminent
  • 366714 03/16/12 15:59 gilman day shift summary
  • 366712 03/16/12 15:54 jmaxwell [Target] New Target Position Offset Method
  • 366704 03/16/12 15:32 jmaxwell [Target] EPICS Reporting Fix
  • 366703 03/16/12 15:27 Fansler LD1 Cryo/ Late Entry
  • 366702 03/16/12 15:18 M. Cummings labeling of plots
  • 366701 03/16/12 15:13 M. Cummings Spot++ Plots for runs 3216/22296
  • 366699 03/16/12 14:56 jmaxwell [Target] Moved Target to Carbon, New TE
  • 366698 03/16/12 14:46 jmaxwell [Target] New Target Position Offset
  • 366697 03/16/12 14:23 M. Cummings Sieve Slits are IN
  • 366692 03/16/12 13:45 yawei Hall C changing half wave plate from out->in
  • 366691 03/16/12 13:43 Chao Scaler ring buffer channel mapping problem
  • 366690 03/16/12 13:13 A. Camsonne Check Right HRS magnet settings
  • 366689 03/16/12 13:08 gilman GenTools for shift checklist
  • 366688 03/16/12 13:03 mhuang RHRS acceptance oddity
  • 366687 03/16/12 12:53 V. Sulkosky Checked Sieve-Slits to Be OUT
  • 366686 03/16/12 12:32 M. Cummings/ RHRS T1/deadtime problems
  • 366685 03/16/12 12:11 brads LHRS dipole field tripped
  • 366684 03/16/12 12:06 gilman chicane magnet
  • 366683 03/16/12 12:03 jmaxwell [Target] New Baselines Taken
  • 366682 03/16/12 11:57 jmaxwell [Target] Entry Finished
  • 366680 03/16/12 11:28 brads GEp/G2p status update and Run Plan sketch -- Friday
  • 366679 03/16/12 11:23 jmaxwell [Target] Opportunistic Entry For NMR Tune, uWave Mixer
  • 366670 03/16/12 11:05 jmaxwell [Target] TE Cup Position
  • 366665 03/16/12 10:56 brads Accelerator down for ~4 hours
  • 366664 03/16/12 10:42 Chao Mapping in db_hel.dat
  • 366663 03/16/12 09:35 D. Keller [Target] TEs and first g2p polarization
  • 366662 03/16/12 09:26 gilman controlled access
  • 366658 03/16/12 08:44 brads Trigger rate anomolies
  • 366653 03/16/12 08:00 pzhu lumi detector
  • 366652 03/16/12 07:56 R. Zielinski Owl Shift Summary
  • 366651 03/16/12 06:53 pzhu slow raster size calibration
  • 366650 03/16/12 06:29 pzhu center beam position check for g2p production run
  • 366649 03/16/12 06:28 R. Zielinski RHRS Scaler/Rate oddity
  • 366630 03/16/12 05:06 R. Zielinski Pengjia trying to center beam on target
  • 366625 03/16/12 04:46 R. Zielinski Hall C changing Half-Wave Plate IHWP to OUT
  • 366615 03/16/12 04:23 pzhu changed the septum to 718.1A
  • 366610 03/16/12 04:08 R. Zielinski MCC sending beam now
  • 366609 03/16/12 04:07 R. Zielinski VDC's ramped back up
  • 366606 03/16/12 04:06 hkkang scaler info 22288(R) 3207(L)
  • 366599 03/16/12 03:35 R. Zielinski Ramping down VDCs as precaution for 1ua beam
  • 366598 03/16/12 02:34 R. Zielinski Still waiting on MCC to tune beam
  • Thursday

  • 366597 03/15/12 23:58 yawei swing shift summary
  • 366596 03/15/12 23:34 yawei Hall C changed the half wave plate
  • 366595 03/15/12 22:52 yawei MCC informing beam off for 20mins
  • 366594 03/15/12 22:35 friedman 90 degrees asymmetry uncertainty
  • 366593 03/15/12 22:07 yawei MCC starting to tune beam!
  • 366592 03/15/12 21:16 yawei Hall A tools
  • 366591 03/15/12 20:48 pzhu BCM calibration for third arm by using one point
  • 366590 03/15/12 20:25 yuxiang [target]
  • 366589 03/15/12 19:21 V. Sulkosky/ Cosmics Runs During Today's Access
  • 366588 03/15/12 19:17 T. Badman [Target] magnet ramped to 2.5T
  • 366587 03/15/12 19:07 yuxiang zhao [target]ramp up target magnet at 7pm
  • 366586 03/15/12 19:04 brads Requested that we go to Power Permit -- we're ready for beam
  • 366585 03/15/12 19:03 jmaxwell [Target] Buffer Dewar Valve Work Finished
  • 366584 03/15/12 18:59 V. Sulkosky/ RHRS Deadtime and DAQ Changes
  • 366583 03/15/12 18:57 yawei Septa fields on
  • 366582 03/15/12 18:35 brads GEp/G2p status update -- Thursday
  • 366581 03/15/12 18:16 pzhu sieve is out
  • 366580 03/15/12 18:00 yawei LHRS VDC HV trip
  • 366579 03/15/12 17:48 pzhu BCM calibration for left and right arm by using one point
  • 366578 03/15/12 17:21 Folts fence /tech
  • 366577 03/15/12 16:44 yawei bring back LHRS Q2
  • 366576 03/15/12 16:25 yawei LHRS Q2 trip
  • 366575 03/15/12 16:07 Fansler Cryo
  • 366574 03/15/12 15:59 gilman day shift summary
  • 366573 03/15/12 15:50 T. Badman start/end of run epics reporting updated
  • 366558 03/15/12 14:23 jmaxwell [Target] Mixer Replaced with Dump
  • 366557 03/15/12 14:07 gilman Q1's trip
  • 366556 03/15/12 13:59 kamyers [Target] magnet level ok
  • 366555 03/15/12 13:48 kamyers [Target] magnet level starting to come down
  • 366554 03/15/12 13:43 Fansler Hall A Q1's
  • 366553 03/15/12 11:41 jmaxwell [Target] uWave Mixer & Valve Actuator Removed
  • 366552 03/15/12 11:17 brads Hot Spot on HA Septum Magnet exit line
  • 366551 03/15/12 11:07 kamyers [Target] Buffer dewer level
  • 366550 03/15/12 10:58 gilman GenTools for shift checklist
  • 366549 03/15/12 10:49 brads Hall Survey from this morning and yesterday
  • 366548 03/15/12 10:27 brads Target group will be replacing encoder on Dewar valve
  • 366547 03/15/12 10:20 folts gate access
  • 366546 03/15/12 09:10 brads Hall moving to controlled access
  • 366545 03/15/12 09:01 kamyers [Target] Magnet ramped to zero
  • 366544 03/15/12 08:43 gilman ramping septa magnets down
  • 366543 03/15/12 08:05 M. Cummings Owl Shift Summary
  • 366542 03/15/12 07:18 M. Cummings Hall C changed IHWP to IN
  • 366541 03/15/12 04:03 M. Cummings Hall C moved IHWP2 (passive) to IN
  • 366540 03/15/12 03:31 D. Keller [Target] supply valve
  • 366539 03/15/12 02:19 M. Cummings MCC still tuning beam
  • 366538 03/15/12 01:52 D. Keller [Target] New Baseline
  • 366537 03/15/12 01:06 M. Cummings Hall C changed IHWP to OUT - alarm worked
  • 366536 03/15/12 00:31 M. Cummings MCC sending tune beam
  • 366531 03/15/12 00:19 M. Cummings Hall A Tools
  • 366526 03/15/12 00:13 gilman no beam - MCC reports FSD/software issues
  • 366521 03/15/12 00:00 gilman swing shift summary
  • Wednesday

  • 366520 03/14/12 23:17 gilman controlled access
  • 366519 03/14/12 23:03 jmaxwell [Target] Buffer Transfer Valve Controlled Access
  • 366518 03/14/12 22:29 gilman gen tools for shift checklist
  • 366517 03/14/12 21:56 T. Badman [Target] TE result correction
  • 366516 03/14/12 21:50 T. Badman [Target] TE results
  • 366515 03/14/12 21:34 Folts VESDA alarm/tech
  • 366514 03/14/12 21:09 Chao 3rd arm helicity signal is back
  • 366507 03/14/12 20:37 A. Camsonne Error in BCM calibration constant for lastr night runs
  • 366506 03/14/12 20:36 gilman controlled access requested
  • 366505 03/14/12 20:25 T. Badman [Target] magnet dewar valve completely open
  • 366500 03/14/12 18:52 Chao 3rd arm Helicity signal do not work
  • 366499 03/14/12 18:34 brads Updates to alarm handler -- new HALFWAVE-PLATE_1
  • 366498 03/14/12 18:24 gilman septa tripped
  • 366497 03/14/12 18:15 gilman iochla crashed again
  • 366496 03/14/12 18:10 gilman septa tripped during tuning
  • 366495 03/14/12 17:55 yawei target->home
  • 366492 03/14/12 17:35 brads Updated firefox profiles on adaq cluster
  • 366491 03/14/12 16:55 gilman mcc needs controlled access
  • 366490 03/14/12 16:36 yawei target magnet ramped up to 2.51T.
  • 366489 03/14/12 16:12 Jixie Zhang Sieve is out
  • 366488 03/14/12 16:10 A. Camsonne BCM receiver slope updated
  • 366487 03/14/12 16:06 Punjabi shift summary
  • 366480 03/14/12 15:40 Jixie Set both HRS to 2.228, septa are already ramped up
  • 366479 03/14/12 13:40 Fansler Checklist
  • 366478 03/14/12 12:54 marble Hall A Venting and Pumpdown Of Beam line between Experiment Runs
  • 366477 03/14/12 12:00 brads G2p/GEp Status update -- Wednesday
  • 366476 03/14/12 11:40 R. Michaels auto ABU to use IBC1H00BCMAVG
  • 366474 03/14/12 09:58 mhuang/chao Straight through optics summary
  • 366473 03/14/12 09:55 Ole Hansen Optics done, asking for access
  • 366469 03/14/12 09:47 Ole Hansen Hall A Tools snapshot
  • 366465 03/14/12 09:43 Chao Scaler screenshot for pointing study
  • 366464 03/14/12 09:37 Ole Hansen BEam stripchart
  • 366461 03/14/12 09:31 Ole Hansen Current stripchart
  • 366460 03/14/12 09:31 Ole Hansen Run 3197 is junk
  • 366455 03/14/12 09:29 Ole Hansen Doing pointing runs now
  • 366451 03/14/12 09:24 Ole Hansen Magnet current stripcharts
  • 366450 03/14/12 09:23 yuxiang zhao [target] fill the LHe into the nose at 9:200
  • 366448 03/14/12 09:23 Ole Hansen hacsbc2 time is off
  • 366445 03/14/12 09:18 Ole Hansen Hall C HWP change
  • 366444 03/14/12 09:17 Ole Hansen spot_L 3rd screen
  • 366443 03/14/12 09:17 Ole Hansen spot_L
  • 366442 03/14/12 09:14 Ole Hansen Septa curret snapshots
  • 366440 03/14/12 09:14 Ole Hansen Hall A tools snapshot
  • 366438 03/14/12 09:13 Ole Hansen Fast raster on
  • 366437 03/14/12 09:08 Ole Hansen LHRS to 0%
  • 366435 03/14/12 09:04 Ole Hansen Run plan
  • 366434 03/14/12 08:59 Ole Hansen Septa stripchart
  • 366433 03/14/12 08:59 mhuang scaler screen 3194 4%
  • 366431 03/14/12 08:55 Ole Hansen Left septum trip
  • 366426 03/14/12 08:41 Ole Hansen Septa current snapshots
  • 366425 03/14/12 08:40 Ole Hansen Hall A tools snapshot
  • 366424 03/14/12 08:36 Ole Hansen LHRS to +4%
  • 366423 03/14/12 08:36 Chao run 22266 dp is 3.5%
  • 366421 03/14/12 08:34 Chao Scaler snapshot LHRS delta 3%
  • 366420 03/14/12 08:30 Ole Hansen Cycle RHRS quads
  • 366419 03/14/12 08:28 Ole Hansen RHRS septum to 725.4A
  • 366417 03/14/12 08:28 Ole Hansen RHRS to 0%
  • 366413 03/14/12 08:24 Ole Hansen Hall A Tools snapshot
  • 366412 03/14/12 08:23 Ole Hansen Septa history
  • 366411 03/14/12 08:23 Ole Hansen Septum screen snapshots +3%
  • 366409 03/14/12 08:19 Ole Hansen Run 3189 is junk
  • 366407 03/14/12 08:16 mhuang scaler snapshot 3% 3189
  • 366405 03/14/12 08:12 Ole Hansen Left HRS to +3%
  • 366403 03/14/12 08:10 gilman shift summary
  • 366400 03/14/12 08:07 mhuang scaler snapshot 3188 2%|22266 3.5%
  • 366395 03/14/12 07:59 gilman spot++
  • 366394 03/14/12 07:52 gilman left arm to 2%.
  • 366388 03/14/12 07:30 Chao Scaler snapshot for left arm delta 0%
  • 366387 03/14/12 07:29 gilman right arm not quite settled
  • 366385 03/14/12 07:27 gilman left arm to 0%
  • 366382 03/14/12 07:23 hkkang RHRS scaler
  • 366379 03/14/12 07:17 gilman right arm to 2%
  • 366376 03/14/12 07:03 mhuang scaler snapshot,3183,-2%
  • 366372 03/14/12 06:56 gilman hrs-left to -2%
  • 366370 03/14/12 06:50 mhuang scaler snapshot, left dp=-3%,3182
  • 366364 03/14/12 06:32 mhuang Because left septum tripped
  • 366362 03/14/12 06:30 jixie left septum tripped! please watch a strip chart!
  • 366360 03/14/12 06:28 gilman right arm at 0%
  • 366359 03/14/12 06:19 adaq beam check for carbon optics
  • 366357 03/14/12 06:17 gilman hrs-left septum had tripped
  • 366353 03/14/12 06:04 mhuang LHRS rates an order lower than RHRS?
  • 366350 03/14/12 05:54 mhuang RHRS shows similar pattern
  • 366348 03/14/12 05:49 gilman left arm to -3%
  • 366346 03/14/12 05:48 mhuang large negative Vz<-->large momentum events
  • 366345 03/14/12 05:47 gilman right arm set to -2%
  • 366344 03/14/12 05:39 gilman left arm changed to -3%
  • 366343 03/14/12 05:38 Chao Scaler snapshot for left arm delta -3%
  • 366341 03/14/12 05:35 Jixie It looks like cosmic rates. The beam might be tripped when taking snapshot.
  • 366334 03/14/12 05:07 pzhu harp scan done
  • 366327 03/14/12 04:56 pzhu rate:position 0 0
  • 366324 03/14/12 04:54 Chao left delta to -4% right delta to -3.5%
  • 366321 03/14/12 04:48 pzhu rate position -4+0
  • 366317 03/14/12 04:37 pzhu rate,+0+4
  • 366314 03/14/12 04:30 mhuang LHRS p0 moving back to -4% 2.341Gev
  • 366313 03/14/12 04:28 Markowitz target nose emptied
  • 366310 03/14/12 04:19 hkkang scaler screenshot
  • 366309 03/14/12 04:18 Min/Chao LHRS 0% sieve pattern
  • 366308 03/14/12 04:16 gilman Hall C changing have wave plate
  • 366307 03/14/12 04:07 mhuang LHRS same setting run from first day optics
  • 366303 03/14/12 04:05 pzhu +4+0 rate, dp=0
  • 366297 03/14/12 03:53 mhuang LHRS weird large momentum events
  • 366282 03/14/12 03:16 Markowitz Moving target to 12C
  • 366281 03/14/12 03:06 jmaxwell [Target] Buffer Dewar Valve Error Access
  • 366280 03/14/12 02:27 A. Camsonne HARP scan 1H05A 5 uA looks good
  • 366279 03/14/12 02:17 Markowitz [Target] Dewar buffer valve stuck open
  • 366278 03/14/12 01:51 T. Badman modified end of run
  • 366277 03/14/12 01:35 mhuang HRS p0 and septa currents change
  • 366276 03/14/12 01:23 gilman shut valve
  • 366271 03/14/12 00:44 friedman gep optics
  • 366266 03/14/12 00:24 A. Camsonne rebooted iochacp2
  • 366262 03/14/12 00:00 kamyers swing shift summary
  • Tuesday

  • 366261 03/13/12 23:57 kamyers Alexandre and Brad on the way in to address issue
  • 366260 03/13/12 23:52 brads Compton electron detector should have been permanently masked?
  • 366259 03/13/12 23:15 kamyers Compton electron detector FSD issue, appears to be IOC problem
  • 366257 03/13/12 22:47 kamyers MCC sending tune beam
  • 366256 03/13/12 22:44 kamyers IHWP is OUT
  • 366255 03/13/12 22:20 A. Camsonne BCM receiver slope updated
  • 366254 03/13/12 22:17 kamyers Problem with power supply card, will not be fixed tonight
  • 366253 03/13/12 22:09 kamyers/kell [Target] target field on Wall reading 1.3 Tesla, but power supply is off
  • 366252 03/13/12 21:54 A. Camsonne Scaler BCM rough calibration right arm
  • 366251 03/13/12 21:53 A. Camsonne Scaler BCM rough calibration right arm
  • 366250 03/13/12 21:19 kamyers septa to 725.4
  • 366249 03/13/12 21:09 kamyers Momentum at 2.251 GeV
  • 366248 03/13/12 20:46 kamyers problem degaussing the FZ magnets, expert on the way in
  • 366247 03/13/12 20:41 kamyers (techs) LN2 level drop and recovery for left Q2
  • 366246 03/13/12 20:19 kamyers Going to beam permit, but may be a while for beam
  • 366245 03/13/12 19:46 jmaxwell [Target] New Shielding for Bunker
  • 366244 03/13/12 19:27 R. Zielinski mysql Run Database for gep/g2p
  • 366243 03/13/12 19:25 mhuang Sieve IN
  • 366242 03/13/12 19:25 Kalyan/Chao Third arm HV moved to hatsv5 slot1
  • 366241 03/13/12 19:02 pzhu attenuator for slow raster ADC changed
  • 366236 03/13/12 18:46 M. Cummings Shift worker instructions for RHRS deadtime issues
  • 366235 03/13/12 18:46 kamyers Hall is now in Controlled Access
  • 366234 03/13/12 18:29 pzhu BPM ADC saturated for yesterday run 3152
  • 366233 03/13/12 18:06 kamyers Hall A General Tools
  • 366232 03/13/12 17:59 Kalyan Beamline HV crate replaced
  • 366231 03/13/12 17:46 kamyers Going to Controlled Access
  • 366230 03/13/12 16:43 S. Sirca Cryo levels in HRSL and HRSR
  • 366229 03/13/12 16:24 Fansler RD1
  • 366228 03/13/12 16:11 Fansler Checklist
  • 366227 03/13/12 16:09 gilman day shift
  • 366226 03/13/12 16:01 jmaxwell [Target] Polarization Loss Due to Magnet Current Ramp Down
  • 366221 03/13/12 15:40 Kalyan Turning off HV GUI to add a test crate
  • 366213 03/13/12 12:31 brads re: changed attenuate for BPM and raster
  • 366212 03/13/12 11:35 pzhu changed attenuate for BPM and raster,connected helicity to ADC
  • 366211 03/13/12 10:53 marble Hall A Venting and Pumpdown Of Beam line between Experiment Runs
  • 366210 03/13/12 10:50 marble Hall A Venting and Pumpdown Of Beam line between Experiment Runs
  • 366200 03/13/12 08:51 jmaxwell [Target] Buffer Dewar Valve Closed
  • 366199 03/13/12 08:35 jmaxwell [Target] Magnet Ramped Down, Leads Locked
  • 366198 03/13/12 08:09 M. Cummings Owl Shift Summary
  • 366187 03/13/12 07:23 adaq ramping down target magnet
  • 366186 03/13/12 07:22 M. Cummings [Target] magnet is ramping down.
  • 366179 03/13/12 07:05 A. Camsonne BCM calibration using one calorimeter point
  • 366178 03/13/12 07:01 M. Cummings [Target] Buffer dewar level still low
  • 366177 03/13/12 06:59 M. Cummings Called for restricted access
  • 366176 03/13/12 06:49 pzhu no beam from 6:20
  • 366175 03/13/12 06:28 M. Cummings Beam down - problems at the injector
  • 366174 03/13/12 06:25 pzhu beam position change
  • 366173 03/13/12 06:18 adaq moved to 12C target
  • 366171 03/13/12 06:09 adaq target buffer dewar alarm level
  • 366160 03/13/12 05:48 adaq target moved to empty
  • 366159 03/13/12 05:38 M. Cummings Bleedthrough test
  • 366155 03/13/12 05:23 M. Cummings Bleedthrough Test
  • 366154 03/13/12 05:22 A. Camsonne Plots tungsten calorimeter
  • 366140 03/13/12 04:55 M. Cummings starting inelastic data
  • 366136 03/13/12 04:50 pzhu hrs rate
  • 366135 03/13/12 04:43 M. Cummings BCM tests done, waiting for target to finish polarizing
  • 366131 03/13/12 04:04 jmaxwell [Target] Polarizing Bottom Cup Positively
  • 366128 03/13/12 03:51 carlin Target work during access
  • 366125 03/13/12 03:46 M. Cummings IHWP change at 2:55 [late entry]
  • 366124 03/13/12 03:39 M. Cummings Sieve slit is OUT
  • 366123 03/13/12 03:31 Chao Charge Asymmetry for run 3137
  • 366122 03/13/12 03:18 M. Cummings Controlled access for sieve slit change
  • 366119 03/13/12 03:16 pzhu raster calibration
  • 366116 03/13/12 03:11 M. Cummings Septum current changed
  • 366111 03/13/12 03:10 jmaxwell [Target] Target Epics Reporting Fixed
  • 366110 03/13/12 03:09 mhuang/jixie Large negative VertexZ events
  • 366105 03/13/12 03:05 M. Cummings L&R HRS dipole settings changed to 2.050
  • 366096 03/13/12 02:59 jmaxwell [Target] Target Motion Problems
  • 366089 03/13/12 02:42 M. Cummings RHRS Deadtime
  • 366086 03/13/12 02:36 pzhu harp scan record
  • 366079 03/13/12 02:24 Jixie scaler rates for 22227|3139
  • 366072 03/13/12 02:06 jmaxwell [Target] Target Magnet Current Loss Again
  • 366069 03/13/12 01:40 jmaxwell [Target] Followup to Computer Unresponsive
  • 366066 03/13/12 01:27 M. Cummings Hall A General Tools
  • 366065 03/13/12 01:26 M. Cummings Spot++ for run 3136
  • 366060 03/13/12 00:58 jmaxwell [Target] NMR Computer Unresponsive
  • 366057 03/13/12 00:47 pengjia Asked for 100na cw beam
  • 366056 03/13/12 00:43 Jixie, Pengj Harp scan start, faster raster off
  • 366055 03/13/12 00:41 Jixie Right dipole is ready for 2.251 GeV
  • 366054 03/13/12 00:35 Chao 7% inefficiency is normal
  • 366053 03/13/12 00:28 Jixie Changed RHRS to delta=0% (2.251), keep LHRS unchanged
  • 366052 03/13/12 00:27 gilman Swing shift summary
  • 366051 03/13/12 00:27 Sasha Moller polarimeter problems
  • 366050 03/13/12 00:27 Chao Thirdarm HV is also tripped
  • Monday

  • 366049 03/12/12 23:49 romanip High Voltage GUI trips
  • 366048 03/12/12 23:44 Pengjia Re:helicity inefficiency?
  • 366047 03/12/12 23:03 Jixie Faster raster off
  • 366046 03/12/12 22:07 jixie HallC changed HWP out
  • 366045 03/12/12 21:56 brads GEp/G2p status update -- Monday
  • 366043 03/12/12 21:37 Jixie, Sasha Prepared to take moller data
  • 366040 03/12/12 21:24 kamyers [Target] Target nose full
  • 366039 03/12/12 21:23 A. Camsonne BCM receiver 800 nA
  • 366038 03/12/12 21:17 A. Camsonne BCM receiver 400 nA
  • 366037 03/12/12 21:11 A. Camsonne BCM receiver 150 nA
  • 366036 03/12/12 21:07 A. Camsonne BCM receiver 100 nA
  • 366035 03/12/12 21:03 A. Camsonne BCM receiver 75 nA
  • 366034 03/12/12 21:01 brads Run plan for Tuesday, Wednesday, Thursday
  • 366033 03/12/12 21:00 A. Camsonne BCM reading 50 nA
  • 366032 03/12/12 20:56 A. Camsonne BCM readout 25 nA
  • 366031 03/12/12 20:56 mhuang Gep optics summary (better format)
  • 366030 03/12/12 20:55 mhuang Gep optics summary
  • 366028 03/12/12 20:55 friedman helicity inefficiency?
  • 366021 03/12/12 20:48 S. Sirca Raw rates during empty cell runs
  • 366020 03/12/12 20:24 mhuang RHRS has higher rate than LHRS (GEP optics runs)
  • 366019 03/12/12 20:24 kamyers [Target] Target moved to home/parked position
  • 366016 03/12/12 20:20 kamyers Scaler rates for bleedthrough measurement
  • 366015 03/12/12 20:19 gilman pedestal bounce
  • 366014 03/12/12 20:14 gilman bleedthrough measurement with bcm
  • 366011 03/12/12 19:58 mhuang scaler screen snapshot
  • 366006 03/12/12 19:35 S. Sirca Snapshot of raw rates during optics runs
  • 366005 03/12/12 19:32 gilman spot check during optics run
  • 365998 03/12/12 19:18 kamyers [Target] Target moved to C12-40
  • 365997 03/12/12 18:57 A. Camsonne Signal seen on HARP 1H05A
  • 365996 03/12/12 18:40 jmaxwell [Target] Moved to Home, Empty Nose
  • 365995 03/12/12 18:21 A. Camsonne turned slow and fast raster off - asked for 8 uA tuned beam
  • 365994 03/12/12 18:18 jmaxwell [Target] TEs Complete Bottom Result
  • 365993 03/12/12 18:16 jmaxwell [Target] TEs Complete
  • 365992 03/12/12 17:07 pzhu changed roc library in right arm happex
  • 365991 03/12/12 16:52 Higinbotham Using HRS for a Bleed-through Measurement
  • 365990 03/12/12 16:38 gilman gen tools for shift checklist
  • 365989 03/12/12 15:57 R. Zielinski Day Shift Summary
  • 365986 03/12/12 15:23 brads Moller is approved -- will start around 5pm
  • 365985 03/12/12 14:37 A. Camsonne Harp preamplifier swapped
  • 365984 03/12/12 13:48 R. Zielinski Sieve Slits IN
  • 365983 03/12/12 13:45 brads Initial plan for when beam returns -- Monday
  • 365982 03/12/12 12:34 duve [target] Turned off microwaves, starting TE
  • 365981 03/12/12 12:28 duve [target] Turned off microwaves, starting TE
  • 365980 03/12/12 12:25 R. Zielinski Requested Controlled Access
  • 365979 03/12/12 12:07 R. Zielinski Beam down for Spot Move
  • 365973 03/12/12 11:48 Ole Hansen Good move
  • 365965 03/12/12 11:29 R. Zielinski Hall C changing Half-Wave Plate
  • 365952 03/12/12 10:09 R. Zielinski Follow Up for Pengjia
  • 365951 03/12/12 09:32 Kalyan Re: sync roc3 and roc5
  • 365950 03/12/12 09:32 R. Zielinski Update on syncROC for 3112
  • 365949 03/12/12 09:32 R. Michaels auto ABU probably meaningless, do it by hand
  • 365948 03/12/12 09:16 yawei stop runs since no beam for next 30 min
  • 365944 03/12/12 09:12 R. Michaels adjust def'n of beam on for auto ABU
  • 365943 03/12/12 09:06 Kalyan Target encoder position in start of run
  • 365935 03/12/12 08:32 R. Zielinski SPOT ++ for LHRS
  • 365934 03/12/12 08:13 R. Zielinski Hall A General Tools Screen
  • 365932 03/12/12 08:08 pzhu moved Rootfiles directory link from /adaql10/work1 to /fs/work1
  • 365929 03/12/12 08:07 friedman owl shift summary
  • 365919 03/12/12 06:43 pzhu right arm happex disconnect problem
  • 365914 03/12/12 06:21 friedman Wrong IHWP entry for runs 3112-3118, 22199-22204 and 40339-40356
  • 365910 03/12/12 05:46 adaq target screens
  • 365902 03/12/12 05:18 friedman IWHP is now IN
  • 365872 03/12/12 03:39 pzhu L_arm s2 compact plot
  • 365867 03/12/12 03:22 pzhu reduced slow raster size from 1.5V to 1.45V
  • 365859 03/12/12 03:12 friedman bleedthrough is zero
  • 365847 03/12/12 02:21 Chao Roughly estimation of the statistical uncertainty of charge asymmetry of run 22194
  • 365846 03/12/12 02:00 pzhu sync roc3 and roc5
  • 365845 03/12/12 01:58 pzhu slow raster ADC saturated
  • 365838 03/12/12 01:45 adaq target screen grab
  • 365835 03/12/12 01:10 friedman IHWP changed - now OUT
  • 365829 03/12/12 00:53 jmaxwell [Target] New Baseline on Top, Basefile Bug, Plan for Night
  • 365828 03/12/12 00:50 friedman hall a tools
  • 365825 03/12/12 00:44 friedman beam back on
  • 365824 03/12/12 00:34 friedman beam stopped
  • 365820 03/12/12 00:27 friedman spot++
  • 365813 03/12/12 00:08 slifer Expected Yields
  • 365812 03/12/12 00:07 A. Camsonne BCM calibration and pedestal
  • 365811 03/12/12 00:05 S. Sirca Swing shift SUMMARY
  • Sunday

  • 365810 03/11/12 23:53 Chao Charge Asymmetry for run 22193
  • 365805 03/11/12 23:05 S. Sirca How to figure out what current we are getting
  • 365794 03/11/12 22:12 jie cold restart for third arm
  • 365791 03/11/12 21:53 slifer shift checklist
  • 365783 03/11/12 21:12 R. Zielinski Passive Half-Wave Plate added to start of run HALOG post
  • 365779 03/11/12 20:43 jie start of run #. epics not exit for third arm
  • 365776 03/11/12 20:37 S. Sirca Laser Controls in GEN_E_TOOLS screen accessible to Hall A?
  • 365770 03/11/12 20:05 moser Hall Bleed Through Measurement
  • 365767 03/11/12 19:56 pzhu normal roc26 process
  • 365761 03/11/12 19:21 jmaxwell [Target] Epics
  • 365760 03/11/12 19:13 slifer Target Vars not updating
  • 365756 03/11/12 18:53 jie correction: 22186 is junk run ( not 21186)
  • 365754 03/11/12 18:50 Jie 21186 is junk run
  • 365753 03/11/12 18:47 pzhu disconnected roc26
  • 365747 03/11/12 18:21 S. Sirca IHWP nomenclature
  • 365746 03/11/12 17:53 jie 3096~3098 are junk runs
  • 365745 03/11/12 17:52 jie cold restarts for LHRS
  • 365736 03/11/12 17:10 S. Sirca Persistent wrong labeling of IHWP status
  • 365730 03/11/12 16:55 brads GEp/G2p status update -- Sunday
  • 365727 03/11/12 16:36 S. Sirca Hall A Tools
  • 365723 03/11/12 16:30 S. Sirca Spot++ for run 3093
  • 365722 03/11/12 16:05 M. Cummings Day Shift Summary
  • 365715 03/11/12 15:47 R. Michaels byte mismathed data files
  • 365708 03/11/12 15:12 josemans Scalers for LHRS & RHRS
  • 365697 03/11/12 13:50 josemans LHRS - #3089
  • 365696 03/11/12 13:48 josemans RHRS - #22178
  • 365695 03/11/12 13:46 A. Camsonne ibcm1 is Hall C current
  • 365688 03/11/12 13:28 M. Cummings BCM A voltage/current dropped to zero
  • 365687 03/11/12 13:13 M. Cummings Should also include 3088
  • 365686 03/11/12 13:11 Chao Update the 3rd arm start-of-run and end-of-run scripts
  • 365685 03/11/12 13:09 M. Cummings Correction: Target type is NH3 Top
  • 365682 03/11/12 13:05 brads re: Bleedthrough test
  • 365680 03/11/12 13:01 A. Camsonne Rough BCM calibration can have offset
  • 365676 03/11/12 12:54 josemans Happex Rebooted 22175
  • 365670 03/11/12 12:35 M. Cummings spot++ for run 3086
  • 365669 03/11/12 12:32 M. Cummings Bleedthrough test
  • 365666 03/11/12 12:24 duve [Target] Target Magnet Current spontaneity
  • 365660 03/11/12 11:41 M. Cummings Beam is back
  • 365657 03/11/12 11:31 M. Cummings Beam Stopped - A2 interception high
  • 365654 03/11/12 11:22 M. Cummings Beam is back
  • 365651 03/11/12 11:06 jmaxwell [Target] Polarizing Top Negatively
  • 365650 03/11/12 10:26 jmaxwell [Target] Anneal Complete
  • 365649 03/11/12 09:00 jmaxwell [Target] Controlled Access for First Anneal
  • 365648 03/11/12 08:51 friedman owl shift summary
  • 365647 03/11/12 08:43 M. Cummings Stop beam for target anneal
  • 365643 03/11/12 08:41 M. Cummings Hall A Tools
  • 365642 03/11/12 08:34 T. Badman wrong prescales for third arm!
  • 365631 03/11/12 07:45 T. Badman slow raster run 3082
  • 365629 03/11/12 07:36 T. Badman ROC26 disconnect again
  • 365628 03/11/12 07:32 T. Badman ROC26 error, rebooted on RHRS run 22168
  • 365619 03/11/12 06:58 friedman Can't read beam position
  • 365615 03/11/12 06:48 friedman IHWP changed - now OUT
  • 365612 03/11/12 06:42 T. Badman ROC26 disconnected
  • 365611 03/11/12 06:39 Higinbotham 100 nA
  • 365610 03/11/12 06:24 friedman spot++
  • 365599 03/11/12 06:08 friedman decrease beam current
  • 365596 03/11/12 06:03 T. Badman beam back on
  • 365593 03/11/12 05:41 friedman beam stop
  • 365590 03/11/12 05:30 slifer Pt
  • 365585 03/11/12 05:01 friedman polarization constantly drops
  • 365584 03/11/12 04:44 friedman dp vs ph by paddles
  • 365579 03/11/12 04:02 T. Badman restarted third arm coda
  • 365571 03/11/12 03:55 T. Badman Slow raster pattern
  • 365570 03/11/12 03:37 friedman bleedthrough
  • 365569 03/11/12 03:23 pzhu happex ADC gain for BCM changed
  • 365568 03/11/12 03:19 T. Badman Coda crashed run 22158
  • 365563 03/11/12 03:04 friedman IHWP changed in middle of runs
  • 365561 03/11/12 03:02 friedman IHWP changed
  • 365555 03/11/12 01:56 T. Badman Third arm start of run
  • 365550 03/11/12 01:33 friedman beam back on
  • 365549 03/11/12 01:30 jmaxwell [Target] Polarizing Negative on Bottom Target
  • 365548 03/11/12 01:22 Kalyan Re: Alarm handler white alarm for two mystery HVs... Problem?
  • 365547 03/11/12 01:10 brads (techs) LN2 level on left arm is flakey
  • 365546 03/11/12 01:03 friedman beam stop
  • 365545 03/11/12 00:51 M. Yurov [target] Change the target
  • 365540 03/11/12 00:41 friedman third arm coda restarted
  • 365535 03/11/12 00:22 friedman spot++
  • 365530 03/11/12 00:10 kamyers Swing shift summary
  • 365529 03/11/12 00:06 Chao Charge asymmetry of run 3065
  • Saturday

  • 365528 03/10/12 23:48 Kalyan Please run Third Arm DAQ tonight
  • 365518 03/10/12 22:56 brads Alarm handler white alarm for two mystery HVs... Problem?
  • 365512 03/10/12 22:01 brads GEp/G2p status update
  • 365509 03/10/12 21:54 D. Keller [Target] Target Magnet Current spontaneity
  • 365504 03/10/12 21:45 kamyers IHWP is now OUT
  • 365501 03/10/12 21:32 kamyers No start run screen/entry for 22148
  • 365500 03/10/12 21:21 D. Keller [Target] Target Magnet Current spontaneity
  • 365499 03/10/12 21:20 kamyers Hall A orbit lock now on
  • 365491 03/10/12 20:45 mhuang update DB 20120306/db_L.vdc.dat
  • 365486 03/10/12 20:24 pzhu right arm happex disconnect sometimes
  • 365485 03/10/12 20:23 mhuang a closer look at CH2 runs
  • 365484 03/10/12 20:22 kamyers Hall A general tools screen
  • 365483 03/10/12 20:20 kamyers large bleedthrough for the following "production" runs
  • 365481 03/10/12 19:56 kamyers slit position should be 14
  • 365479 03/10/12 19:54 jmaxwell [Target] Slow Raster still looks good
  • 365478 03/10/12 19:52 kamyers Bleedthrough significantly reduced
  • 365477 03/10/12 19:48 jie there is significant hall c bleedthrough
  • 365465 03/10/12 19:24 mhuang fp_x plots for refercence to dp plots
  • 365462 03/10/12 18:58 jmaxwell [Target] Magnetic Field Moved Again
  • 365459 03/10/12 18:53 jmaxwell [Target] Slow Raster Plot
  • 365456 03/10/12 18:49 kamyers bcm readings with 100 nA
  • 365455 03/10/12 18:47 kamyers run 3058 production, IHWP is IN
  • 365454 03/10/12 18:46 kamyers this run junk (3057)
  • 365449 03/10/12 18:35 kamyers spot L plots run 3056
  • 365448 03/10/12 18:32 kamyers small amount of leakage beam
  • 365440 03/10/12 18:10 jmaxwell [Target] Run Valve Back on PID
  • 365439 03/10/12 17:59 kamyers fast raster is set to 2x2, slow raster is on at 1.5V
  • 365437 03/10/12 17:55 kamyers IHWP is IN
  • 365436 03/10/12 17:54 jmaxwell Slow Raster On
  • 365434 03/10/12 17:46 jmaxwell [Target] TE Measurements
  • 365433 03/10/12 17:41 jmaxwell [Target] Polariing Top, Positive
  • 365430 03/10/12 16:54 M. Cummings adjusted HV settings
  • 365429 03/10/12 16:51 kamyers septum currents at 718.1 A
  • 365425 03/10/12 16:30 kamyers momentum set to 2.228 GeV in both arms
  • 365424 03/10/12 16:26 Chao Sieve OUT
  • 365423 03/10/12 16:18 R. Zielinski Enabled EDTM on both arms
  • 365422 03/10/12 16:09 Chao set HRS momentum to 2.228GeV
  • 365421 03/10/12 16:06 Chao Day Shift Summary
  • 365420 03/10/12 15:55 Chao Hall A General Tools
  • 365419 03/10/12 14:43 A. Camsonne Rough BCM calibration
  • 365418 03/10/12 14:39 Chao IHWP Change
  • 365417 03/10/12 14:35 Chao/Yunchen Comment for run 40296
  • 365416 03/10/12 14:27 A. Camsonne Runs 3053 and 22133 BCM pedestal run
  • 365412 03/10/12 14:08 A. Camsonne Requested beam off for TE measurement
  • 365406 03/10/12 14:00 A. Camsonne BCM reading Hall A slit closed
  • 365405 03/10/12 13:57 jmaxwell [Target] Move to Top for TE
  • 365404 03/10/12 13:56 Chao vertex z distribution for left arm run 3041
  • 365401 03/10/12 13:53 A. Camsonne BCM screen withh 100 nA
  • 365400 03/10/12 13:49 mhuang Gep dp spectrum plot RHRS #22120
  • 365399 03/10/12 13:47 Chao Sieve pattern for right arm run 22120
  • 365396 03/10/12 13:35 A. Camsonne BCM receiver gain increased for low current running
  • 365395 03/10/12 13:04 A. Camsonne Old BCM receiver gains settings
  • 365394 03/10/12 12:52 Jack(Techs) Right Dipole behaving normally
  • 365393 03/10/12 12:43 jpierce [target] Moved target out
  • 365392 03/10/12 12:24 jmaxwell [Target] Moved to Top
  • 365391 03/10/12 12:17 jmaxwell [Target] Taking Baselines
  • 365390 03/10/12 12:11 jmaxwell [Target] Nose PID loop back on
  • 365389 03/10/12 12:07 jmaxwell [Target] Moved to Bottom
  • 365388 03/10/12 11:55 jmaxwell [Target] Moved to Top, Filling Nose
  • 365387 03/10/12 10:52 Jack(Techs) Right Dipole ramping down
  • 365386 03/10/12 09:28 Jixie Controlled access
  • 365385 03/10/12 09:05 Jixie MCC took beam away
  • 365384 03/10/12 09:05 Jixie Moved beam to (0,-4), Not settled
  • 365383 03/10/12 08:42 Yuncheng Run 3050, Beam Position (0,4)
  • 365382 03/10/12 08:39 Yuncheng Run 22129, Target, Beam Position (0,4)
  • 365379 03/10/12 08:34 Yuncheng Run 22129 FP, Beam Position (0,4)
  • 365378 03/10/12 08:31 Yuncheng Run 3050, beam Position (0,4)
  • 365377 03/10/12 08:26 Yuncheng Run 3051, beam position (0,4)
  • 365374 03/10/12 08:23 Jixie Beam is back
  • 365373 03/10/12 08:15 Jixie Beam is not available
  • 365370 03/10/12 08:10 yuncheng replay 3050
  • 365369 03/10/12 08:03 Jixie Move beam to (0,4)
  • 365368 03/10/12 08:01 pzhu owl shift summary
  • 365367 03/10/12 08:01 Jixie Hall C changed HWP
  • 365366 03/10/12 08:00 efuchey Scalers runs 3050/22129
  • 365357 03/10/12 07:38 efuchey Scalers runs 3046/22125
  • 365354 03/10/12 07:36 Jixie Moved beam to (4,4)
  • 365343 03/10/12 07:13 Jixie Moved beam to (-4,0)
  • 365338 03/10/12 07:02 Jixie Hall C changed HWP
  • 365333 03/10/12 06:46 Jixie Move beam to (4,0)
  • 365332 03/10/12 06:43 efuchey scalers for runs 3043/22122
  • 365323 03/10/12 06:14 Jixie Beam is at (0,0)
  • 365322 03/10/12 06:12 efuchey scalers runs 3041/22120
  • 365319 03/10/12 06:09 Jixie optics restarted! Move beam to (0,0)
  • 365318 03/10/12 06:02 Jixie Hall C change HWP
  • 365317 03/10/12 06:02 Jixie Wrong momentum control for both arm from Swing shift
  • 365313 03/10/12 05:47 efuchey Scalers runs 3040/22119
  • 365300 03/10/12 05:19 efuchey scalers, run 3035/22115
  • 365299 03/10/12 05:18 efuchey scalers run 3035/22115
  • 365290 03/10/12 04:58 jixie Hall C changed HWP
  • 365289 03/10/12 04:55 efuchey scalers runs 3033/22113
  • 365288 03/10/12 04:53 M. Yurov [taget] magnet warm return flow
  • 365285 03/10/12 04:51 Jixie Asked MCC to move beam to (-4,-4)
  • 365277 03/10/12 04:29 efuchey roc 26 crashed
  • 365276 03/10/12 04:26 efuchey scalers run 3031/22110
  • 365275 03/10/12 04:23 Jixie (0,-6) is too low, moved beam to (0.-4)
  • 365268 03/10/12 04:12 Jixie Move beam to (0,-12)
  • 365261 03/10/12 03:54 efuchey scalers: beam pos scan (0,6)
  • 365260 03/10/12 03:54 Jixie Hall C change HWP
  • 365259 03/10/12 03:52 Jixie Beam finally reached (0,6)
  • 365256 03/10/12 03:36 pzhu slow raster size for new configuration
  • 365255 03/10/12 03:32 Jixie Finished (0,0) data taking, move beam to (0,6)
  • 365252 03/10/12 03:19 Jixie Start optics beam position scan
  • 365245 03/10/12 03:08 Jixie Turn off the slow raster
  • 365244 03/10/12 03:07 efuchey spot L and R for runs 3025/22104
  • 365241 03/10/12 03:05 Jixie Slow raster on
  • 365239 03/10/12 03:04 Jixie changed beam position
  • 365233 03/10/12 02:54 Jixie Slow raster off
  • 365232 03/10/12 02:54 Jixie Hall C change HWP
  • 365229 03/10/12 02:51 efuchey run 22100 right arm
  • 365226 03/10/12 02:43 efuchey run 3021 replay
  • 365223 03/10/12 02:39 pzhu slow raster shot
  • 365222 03/10/12 02:39 Jixie Slow raser on, change BPMB x from -1.7 to -1.0
  • 365221 03/10/12 02:35 efuchey spot L and R w/ slow raster
  • 365218 03/10/12 02:33 M. Yurov [target] Magnet current has decreased
  • 365215 03/10/12 02:26 Jixie Tools window
  • 365214 03/10/12 02:24 efuchey scalers
  • 365213 03/10/12 02:22 M.Yurov [target] Top/Bottom temp. camera is off
  • 365208 03/10/12 02:16 efuchey scalers for delta_L 0, delta_R +1
  • 365207 03/10/12 02:13 efuchey spot L 3019
  • 365202 03/10/12 02:03 jmaxwell [Target] No Helium Running
  • 365199 03/10/12 01:37 pzhu start optics data
  • 365198 03/10/12 01:37 M. Yurov [target] Runing with low helium
  • 365193 03/10/12 01:26 M. [target] Moved target
  • 365192 03/10/12 01:08 adaq no signal for 1H05A harp
  • 365191 03/10/12 01:03 M. Yurov [target] Magnet current has decreased (same as Entry number 365126/365189)
  • 365190 03/10/12 00:55 M Yurov [target] Started to drain helium from the nose
  • 365189 03/10/12 00:51 M. Yurov [target] Magnet current has decreased (same as Entry number 365126)
  • 365188 03/10/12 00:40 Kalyan Re: fast raster calibration failed
  • 365179 03/10/12 00:26 pzhu 100nA CW1H04A harpscan
  • 365174 03/10/12 00:13 jmaxwell [Target] Slow Raster Size Calibration
  • Friday

  • 365172 03/09/12 23:58 kamyers Swing shift summary
  • 365171 03/09/12 23:49 jpierce Magnet Fill
  • 365170 03/09/12 23:49 J. Shields [Target] moved target home
  • 365169 03/09/12 23:46 pzhu fast raster calibration failed
  • 365168 03/09/12 23:41 kamyers changes to right Q1 and right septum
  • 365164 03/09/12 23:27 kamyers fast raster to 7x7
  • 365163 03/09/12 23:25 pzhu slow raster size calibration
  • 365158 03/09/12 23:12 Kalyan New beam positions for 5T field, 6deg target rotation
  • 365157 03/09/12 23:10 kamyers slow raster off, fast raster set to 10x10 mm2
  • 365156 03/09/12 23:05 kamyers final beam positions for the 5T setting
  • 365155 03/09/12 23:03 pzhu new beam center check:A(-0.1,2.0),B(-1.6,0)
  • 365153 03/09/12 23:01 kamyers beam positions for run 3009
  • 365149 03/09/12 22:53 kamyers beam positions for runb 3008
  • 365144 03/09/12 22:30 J. Shields [Target] moved target to C12 hole
  • 365143 03/09/12 22:25 kamyers beam positions for run 3005
  • 365140 03/09/12 22:14 kamyers beam positions for run 3004
  • 365137 03/09/12 21:59 kamyers beam positions for this run
  • 365132 03/09/12 21:42 J. Shields [Target] Moved target halfway between small holes
  • 365131 03/09/12 21:13 kamyers fast raster off, slow raster on
  • 365130 03/09/12 21:10 J. Shields [Target] moved target to C12-40
  • 365129 03/09/12 20:54 D. Keller [Target] sys normal
  • 365128 03/09/12 20:31 J. Shields [Target] Noticed that the warm return flow (FL91115) suddenly went off scale high (> 200) just before the magnet acted up
  • 365127 03/09/12 20:25 Tech on call Right dipole power supply
  • 03/09/12 03/09/12 20:25 J. Shields [Target] Magnet seems to have decreased current autonomously by 0.5 A. We manually drove it back to 121.825 A.
  • 365125 03/09/12 20:19 kamyers right arm to 2.22799 GeV for tonight
  • 365124 03/09/12 19:59 J. Shields [Target] Nose level reached 50%, self-recovered
  • 365123 03/09/12 19:50 kamyers MCC going to start tuning
  • 365122 03/09/12 19:47 D. Keller [Target] Target Magnet Ramped to 5T
  • 365121 03/09/12 19:39 A. Camsonne Right arm dipole power supply stuck
  • 365120 03/09/12 19:17 kamyers Left and Right quads cycled, LHRS set to 2.251 GeV, RHRS dipole current limited
  • 365119 03/09/12 18:25 kamyers Septum ramped to 725.4 A
  • 365118 03/09/12 18:15 Jack(Techs) UPS on Left Arm replaced
  • 365117 03/09/12 18:09 mhuang/chao Sieve IN
  • 365116 03/09/12 18:03 kamyers Sweep of Hall finished
  • 365115 03/09/12 17:49 M. Cummings T2 rate problem
  • 365114 03/09/12 17:34 pzhu bpm pedestal run
  • 365109 03/09/12 17:32 Jack(Techs) Right Q3 current limit is now 1400 amps
  • 365105 03/09/12 17:29 pzhu bpm pedestal run
  • 365103 03/09/12 17:06 kamyers Work in hall completed, going to start sweep
  • 365092 03/09/12 16:02 Jie day shift summary
  • 365091 03/09/12 15:58 Fansler Checklist
  • 365089 03/09/12 15:51 D. Keller [Target] Target Magnet Ramped Down
  • 365085 03/09/12 15:36 pzhu BPM receiver setting backup
  • 365083 03/09/12 14:40 jmaxwell [Target] Magnet Ramped to 5T
  • 365081 03/09/12 13:38 jmaxwell [Target] Magnet Ramped up to 5T, Now Ramping Down
  • 365076 03/09/12 11:24 brads Beamline <-> Target flange alignment snapshots
  • 365070 03/09/12 10:53 Zhihong Ye Chicane Move completed
  • 365066 03/09/12 09:38 Zhihong Ye G2p Status Update
  • 365063 03/09/12 09:14 M. Cummings/ T2 rate problem
  • 365059 03/09/12 08:18 X. Zheng Restricted Access
  • 365058 03/09/12 08:07 Chao Owl Shift Summary
  • 365057 03/09/12 07:26 Chao Ramped the septa off
  • 365056 03/09/12 07:20 Chao Hall A General Tools
  • 365055 03/09/12 07:12 Chao Beam is down, asked for a full survey of the hall
  • 365050 03/09/12 07:06 R. Zielinski [ Target ] Moved to bottom NH3
  • 365044 03/09/12 06:58 R. Zielinski [ Target ] Moved to Cross-Hair target
  • 365039 03/09/12 06:44 R. Zielinski LHRS Dilution Study Run 2994 Scaler
  • 365034 03/09/12 06:38 Chao Set both arms to P0=2.228GeV
  • 365031 03/09/12 06:32 Chao change IHWP
  • 365030 03/09/12 06:32 R. Zielinski [ Target ]Moved target to TOP NH3
  • 365026 03/09/12 06:23 Chao Spot++ plots for run 2990
  • 365022 03/09/12 06:21 Chao turned slow raster on
  • 365017 03/09/12 06:05 Chao Spot++ plots for run 2989
  • 365006 03/09/12 05:24 Chao called MCC to turn on fast raster
  • 364992 03/09/12 04:41 Chao Correction: It should be moved sieve OUT here
  • 364991 03/09/12 04:36 Chao moved sieve in
  • 364990 03/09/12 04:17 R. Zielinski [ Target ] Refilling nose with helium
  • 364989 03/09/12 04:16 Chao asked MCC for an escorted access
  • 364986 03/09/12 04:09 R. Zielinski LHRS 4% Run 2982 Plots
  • 364984 03/09/12 04:07 Chao set left delta to 4%
  • 364983 03/09/12 04:06 Chao finished delta scan on right arm
  • 364981 03/09/12 04:05 R. Zielinski LHRS 4% Scaler 2982
  • 364978 03/09/12 03:59 Chao set left delta to 3% right delta to 3.5%
  • 364975 03/09/12 03:59 R. Zielinski RHRS 3.5% Run 22058 Plots
  • 364971 03/09/12 03:55 R. Zielinski RHRS 3.5% Scaler 22058
  • 364970 03/09/12 03:54 R. Zielinski LHRS 3% 2981 Plots
  • 364965 03/09/12 03:49 R. Zielinski LHRS 3% Scalers 2980
  • 364957 03/09/12 03:40 R. Zielinski LHRS 2% Run 2978 Plots
  • 364953 03/09/12 03:37 R. Zielinski LHRS Scaler 2% 2978
  • 364951 03/09/12 03:34 Chao Set left delta to 2%
  • 364948 03/09/12 03:26 Chao IHWP change
  • 364947 03/09/12 03:25 R. Zielinski Both DAQs not getting any rate so I rebooted them
  • 364944 03/09/12 03:22 R. Zielinski RHRS 2% 22056 Plots
  • 364940 03/09/12 03:19 Chao set right delta to 2%
  • 364937 03/09/12 03:16 R. Zielinski LHRS 1% Run 2974 Plots
  • 364933 03/09/12 03:12 Chao set left delta to 1%
  • 364932 03/09/12 03:11 R. Zielinski LHRS 1% Scaler 2974
  • 364929 03/09/12 03:01 R. Zielinski LHRS -1% Run 2973 Scalers
  • 364928 03/09/12 03:00 R. Zielinski LHRS -1% Run 2972 Plots
  • 364926 03/09/12 02:55 Chao set left delta to -1%
  • 364922 03/09/12 02:44 R. Zielinski LHRS -2% 2970 Plots
  • 364917 03/09/12 02:39 Chao set right delta to -1%
  • 364916 03/09/12 02:39 R. Zielinski RHRS -1% 22054 Plots
  • 364914 03/09/12 02:34 R. Zielinski RHRS 22054, LHRS 2970 Scaler
  • 364913 03/09/12 02:34 Chao set left delta to -2%
  • 364910 03/09/12 02:24 Chao IHWP Change
  • 364909 03/09/12 02:22 R. Zielinski LHRS -3% 2968 Plots
  • 364906 03/09/12 02:19 R.Zielinski LHRS Scaler-3% 2968
  • 364904 03/09/12 02:13 Chao set left delta to -3%
  • 364903 03/09/12 02:10 R. Zielinski LHRS -4% 2966 Plots
  • 364900 03/09/12 02:03 jmaxwell [Target] Target Position Test Results
  • 364899 03/09/12 02:01 R. Zielinski RHRS -3.5% 22052 Plots
  • 364894 03/09/12 01:51 R. Zielinski LHRS Scaler 2966 -4%
  • 364892 03/09/12 01:48 Chao set left delta to -4%
  • 364887 03/09/12 01:39 Chao move on to optics data taking
  • 364886 03/09/12 01:39 R. Zielinski RHRS 22051 Scaler (-3.5%)
  • 364884 03/09/12 01:36 jmaxwell [Target] Moved to Carbon 40 mil
  • 364883 03/09/12 01:35 Chao Turned slow raster off
  • 364870 03/09/12 00:59 mhuang Sieve patterns as vertical beam position change
  • 364869 03/09/12 00:48 Chao Hall A General Tools
  • 364868 03/09/12 00:32 mhuang Update of dp plot #2921
  • 364867 03/09/12 00:28 mhuang tg_th and dp plots: target OFF vs. target field ON
  • 364866 03/09/12 00:22 jpchen/Jixie Swing shift summary
  • 364865 03/09/12 00:22 jpchen IHWP change
  • 364863 03/09/12 00:01 jpchen Target moved to no target for beam tuning
  • 364862 03/09/12 00:01 jmaxwell [Target] Move to Home
  • Thursday

  • 364856 03/08/12 23:47 pzhu changed slow raster size from 1.4V to 1.5V
  • 364852 03/08/12 23:40 Jixie, JP, J change left delta to 0% for target position test
  • 364850 03/08/12 23:34 jmaxwell [Target] Position Check, Slow Raster 1.4V
  • 364849 03/08/12 23:33 Jixie, JP, J Start taking data, right dipole not ready yet
  • 364848 03/08/12 23:27 Jixie, JP Set left delta to -4% and right hrs delta to -3.5%
  • 364847 03/08/12 23:24 jpchen A short summary
  • 364846 03/08/12 23:08 Jixie,JP Start delta scan. Set delta to 1%,
  • 364845 03/08/12 23:04 yuxiang zhao slow raster on, after beam 1mm(mcc unit) up,run 2953
  • 364844 03/08/12 22:59 J. Shields [Target] optical 12386 linear 3054
  • 364843 03/08/12 22:59 JP, Jixie Correct: Slow Raster is 500mv, BPM position are
  • 364839 03/08/12 22:57 yuxiang zhao R-run 22049 can not be used
  • 364838 03/08/12 22:55 JP, Jixie Correct: Slow Raster is 500mv, BPM position are
  • 364833 03/08/12 22:46 Jixie,JP Moved beam 1 mm up (MCC unit)
  • 364832 03/08/12 22:40 yuxiang zhao slow raster keep changing run 2951
  • 364831 03/08/12 22:39 yuxiang zhao run 2951 slow raster(500mv )on keep changing
  • 364824 03/08/12 22:24 Jixie, JP Turned slow raster on, 500mv
  • 364821 03/08/12 22:18 yuxiang acceptance study I faster raster on run 2950
  • 364820 03/08/12 22:16 yuxiang zhao acceptance study I run 2949
  • 364815 03/08/12 22:12 yuxiang zhao run 2949 acceptance study 1 faster raster on
  • 364814 03/08/12 22:06 JP, Jixie Started optics acceptance calibration
  • 364805 03/08/12 21:44 yuxiang zhao L-run 2947 R-run 22043
  • 364804 03/08/12 21:42 yuxiang zhao replay R-run 22042
  • 364803 03/08/12 21:41 yuxiang zhao replay for Lrun2946
  • 364800 03/08/12 21:35 Jixie Moved beam to (6,0)
  • 364797 03/08/12 21:30 Jixie Title is wrong. it is (0,-4)
  • 364796 03/08/12 21:25 yuxiang zhao Lrun 2946 Rrun 22042
  • 364791 03/08/12 21:14 yuxiang zhao replay for L-run 2944
  • 364790 03/08/12 21:13 yuxiang zhao replay for R-run 22040
  • 364789 03/08/12 21:11 jpchen IHWP change
  • 364784 03/08/12 20:56 yuxiang zhao Lrun2944 R-run 22040
  • 364781 03/08/12 20:55 JP, Jixie Move beam to (-6,0)
  • 364778 03/08/12 20:48 yuxiang zhao replay R-run22038
  • 364777 03/08/12 20:47 yuxiang zhao replay L-run 2942
  • 364774 03/08/12 20:42 J. Shields [Target] Left rate increased to 1.4 kHz. Appears to be an expected result of beam movement.
  • 364771 03/08/12 20:39 Jixie, JP Change beam to (-4,-4)
  • 364770 03/08/12 20:35 yuxiang L-run 2942 R-run 22038
  • 364761 03/08/12 20:14 yuxiang zhao replay R-run 22035
  • 364760 03/08/12 20:13 yuxiang zhao replay L-run 2939
  • 364758 03/08/12 20:10 pzhu slow raster function generator update
  • 364753 03/08/12 20:02 Jixie, JP Move beam to (0,-6)
  • 364747 03/08/12 19:47 yuxiang zhao replay for L-run 2938
  • 364746 03/08/12 19:44 yuxiang zhao replay for R-run 22033
  • 364741 03/08/12 19:37 yuxiang zhao replay for R-run 22032
  • 364740 03/08/12 19:36 yuxiang zhao replay for L-run 2936
  • 364739 03/08/12 19:33 yuxiang zhao L-run 2937 R-run 22033
  • 364736 03/08/12 19:25 Jixie, JP Change beam to (0,-10)
  • 364733 03/08/12 19:12 jpchen Beam IHWP change
  • 364728 03/08/12 18:54 Jixie, JP Beam position is at (0,6)
  • 364724 03/08/12 18:49 Jixie, JP Move beam to (0,8), then change to (0,6)
  • 364718 03/08/12 18:40 yuxiang zhao position scan(0,6)
  • 364716 03/08/12 18:38 J. Shields [Target] Rate momentarily returned to 5 kHz, probably due to beam crash
  • 364713 03/08/12 18:25 jmaxwell [Target] Running with NO Helium
  • 364710 03/08/12 18:23 yuxiang zhao replay R-run 22026
  • 364709 03/08/12 18:19 yuxiang zhao replay L-run 2930
  • 364708 03/08/12 18:18 Jixie, JP Tools snapshot, beam position(0,4)
  • 364707 03/08/12 18:16 jpchen A beam is going through Compton Chicane
  • 364702 03/08/12 18:13 jpchen beam IHWP changed
  • 364701 03/08/12 18:08 Jixie Change beam 4mm vertical up
  • 364700 03/08/12 18:05 yuxiang zhao R-run 2930 L-run 22026
  • 364696 03/08/12 17:58 yuxiang zhao replay for R-run 22024
  • 364695 03/08/12 17:56 yuxiang zhao replay L-run 22024
  • 364694 03/08/12 17:52 yuxiang position scan (0,0) L-run 2929 R-run 22025
  • 364689 03/08/12 17:44 jmaxwell [Target] New Target Positions Updated on Table, in PDP
  • 364688 03/08/12 17:37 yuxiang zhao replay for RHRS run 22023
  • 364687 03/08/12 17:32 yuxiang L-run number 2928 R-run number 22024
  • 364685 03/08/12 17:29 yuxiang replay for LHRS run 2927
  • 364683 03/08/12 17:27 mhuang Adding the sieve pattern plot of LHRS
  • 364680 03/08/12 17:19 yuxiang zhao L-run number 2927 R-run number 22023
  • 364679 03/08/12 17:10 YUXIANG L-run number 2927 R-run number 22023
  • 364678 03/08/12 17:08 R. Michaels revival automated ABU (check it)
  • 364662 03/08/12 16:11 X. Zheng Day shift summary
  • 364661 03/08/12 16:01 mhuang/jixie Target field ON runs: sieve pattern & dp spectrum
  • 364656 03/08/12 15:49 jmaxwell [Target] Closed Run, Turned on uWaves
  • 364646 03/08/12 15:24 X. Zheng Beam nominal position for centering on the target
  • 364645 03/08/12 15:23 X. Zheng Beam nominal position for centering on the target
  • 364644 03/08/12 15:22 kamyers [Target] red alarms on 4K supply temperature
  • 364636 03/08/12 15:09 Jixie Start to take optics data
  • 364630 03/08/12 14:54 X. Zheng Hall C change beam HWP again
  • 364628 03/08/12 14:41 jmaxwell [Target] Position Offset Value
  • 364627 03/08/12 14:36 jmaxwell [Target] Moved to C12 40 mil
  • 364626 03/08/12 14:33 kalyan Slow raster off in prepration for optics runs
  • 364625 03/08/12 14:27 Kalyan BPM strip chart for target center
  • 364614 03/08/12 13:56 Solvignon IHWP change
  • 364611 03/08/12 13:54 kamyers [Target] Centered on small C12 hole target
  • 364608 03/08/12 13:52 jmaxwell [Target] New Position Encoder Offsets (plot)
  • 364583 03/08/12 13:17 X. Zheng slow raster pattern
  • 364581 03/08/12 13:15 Solvignon Reboot ROC26
  • 364576 03/08/12 12:17 X. Zheng Or not!
  • 364575 03/08/12 12:09 X. Zheng Beam back!
  • 364574 03/08/12 10:23 X. Zheng Hall C change beam HWP
  • 364573 03/08/12 10:23 X. Zheng Beam will be down for about 30 minutes
  • 364572 03/08/12 10:21 R. Michaels run file byte mismatch for run 21580
  • 364561 03/08/12 10:00 X. Zheng Beam back
  • 364560 03/08/12 09:59 jmaxwell [Target] Moved to Hole Target
  • 364559 03/08/12 09:57 jmaxwell Slow Raster On
  • 364558 03/08/12 09:50 Zhihong Ye G2p Status Update: Smooth
  • 364557 03/08/12 09:39 X. Zheng Beam delivery problem
  • 364556 03/08/12 09:28 jmaxwell [Target] Moved to Home
  • 364555 03/08/12 09:11 pzhu slow raster size check and harp scan
  • 364554 03/08/12 09:05 Jixie Set septa currents to 725.4A
  • 364553 03/08/12 09:05 jmaxwell [Target] Moved to Carbon Target
  • 364552 03/08/12 08:39 Jixie, Chao Escorted acess to move sieve in
  • 364551 03/08/12 08:34 T. Badman raster size calibration
  • 364550 03/08/12 08:33 X. Zheng Escorted access for inserting the sieve slits
  • 364549 03/08/12 08:29 X. Zheng Hall A Tools and Septum Screen
  • 364548 03/08/12 08:08 M. Cummings Owl Shift Summary
  • 364547 03/08/12 08:07 M. Cummings Cycling LHRS momentum back to 2.251
  • 364546 03/08/12 07:58 M. Cummings Hall C changing half wave plate
  • 364541 03/08/12 07:53 pzhu slow raster size
  • 364536 03/08/12 07:46 marcoapc Scaler - fast raster off, target at 4030526 - runs #21999 (right) and #2902 (left)
  • 364519 03/08/12 07:16 marcoapc Scaler - target at 4172999 - runs #21995 (right) and #2898 (left)
  • 364518 03/08/12 06:51 A. Camsonne Slow raster rough calibration
  • 364511 03/08/12 06:21 M. Cummings LHRS momentum moved to 2.0934
  • 364510 03/08/12 06:20 marcoapc Scaler - target crosshair (1801499) - runs #21993 (right) and #2896 (left)
  • 364503 03/08/12 06:16 marcoapc Scaler - crosshair target (2217749) - runs #21992 (right) and #2895 (left)
  • 364502 03/08/12 06:06 T raster calibration update
  • 364497 03/08/12 05:59 marcoapc Scaler - target at 4000499 - runs #21991 (right) and #2894 (left)
  • 364496 03/08/12 05:55 M. Cummings Hall C changing HWP
  • 364493 03/08/12 05:53 marcoapc Scaler - ps=1 (right) and ps3=1 (left) and ps8=0 (both) - runs #21990 (right) and #2893 (left
  • 364487 03/08/12 05:48 marcoapc Scaler - using ps8=1 and ps1=ps3=0 - runs #21989 (right) and #2892 (left)
  • 364483 03/08/12 05:44 marcoapc Scaler - beam current 50nA CW - runs #21988 (right) and #2891 (left)
  • 364476 03/08/12 05:41 marcoapc Scaler - target at CH2 (3883499) - runs #21987 (right) and #2890 (left)
  • 364472 03/08/12 05:34 marcoapc Scaler - beam position +1,-3 and target at 2217749 - runs #2889 (left) and #21986 (right)
  • 364462 03/08/12 05:18 marcoapc Scaler - target at 1745999
  • 364461 03/08/12 05:17 pzhu slow raster size check
  • 364458 03/08/12 05:16 marcoapc Scaler - target at 1805999 - runs #21983 (right) and #2886 (left)
  • 364453 03/08/12 05:11 marcoapc Scaler - target at crosshair - runs #21982 (right) and #2885 (left)
  • 364450 03/08/12 05:07 adaq [Target] Moved to crosshair
  • 364445 03/08/12 05:04 marcoapc Scaler - target moved to 3765369
  • 364440 03/08/12 05:00 adaq [Target] moved target position to CH2
  • 364439 03/08/12 05:00 marcoapc Scaler - target moved to 38834
  • 364434 03/08/12 04:56 marcoapc Scaler - target moved to 4292609 - runs #21979 (right) and #2882 (left)
  • 364433 03/08/12 04:54 M. Cummings Hall C changing HWP
  • 364428 03/08/12 04:53 marcoapc Scaler - target moved to 4233554 - runs #21978 (rigth) and #2881 (left)
  • 364423 03/08/12 04:49 marcoapc Scaler - target moved position 4174499 - runs #21977 (right) and #2880 (left)
  • 364422 03/08/12 04:44 pzhu raster size check
  • 364419 03/08/12 04:38 adaq [Target] small target movement
  • 364418 03/08/12 04:38 marcoapc Scaler - target moved 2.7mm, up
  • 364410 03/08/12 04:28 pzhu beam position -1-4 to +1-4
  • 364406 03/08/12 04:21 marcoapc Scaler - beam position +1,-4 - runs #21974 (right) and #2877 (left)
  • 364400 03/08/12 04:07 marcoapc Scaler - 500 nA - runs #21973 (right) and #2876 (left)
  • 364393 03/08/12 04:03 marcoapc Scalers - slow raster 1.4V, 200nA - runs #2874 (left) and #21971 (right)
  • 364384 03/08/12 03:55 marcoapc Scaler - raster 2x2, slow raster 1V
  • 364381 03/08/12 03:53 M. Cummings Hall C changing HWP
  • 364377 03/08/12 03:50 marcoapc Scaler - 100nA - runs #2871 (left) and #21968 (right)
  • 364371 03/08/12 03:45 marcoapc Scaler - raster calibration, pulse 50nA, 2x2 fast raster and 500mV slow raster - runs #2870 (left) and #21967 (right)
  • 364370 03/08/12 03:39 pzhu slow raster size calibration
  • 364369 03/08/12 03:39 T. Badman [Target] Moved target for slow raster calibration
  • 364368 03/08/12 03:39 marcoapc Scalers - beam off - preparing BPM + slow raster calib
  • 364367 03/08/12 03:37 M. Cummings Septa turned on
  • 364366 03/08/12 03:35 M. Cummings Harp scans finished
  • 364363 03/08/12 03:25 pzhu harp scan record
  • 364359 03/08/12 02:49 M. Cummings Hall C changing HWP
  • 364351 03/08/12 02:01 M. Cummings Hall A Tools
  • 364346 03/08/12 01:39 M. Cummings Hall C changing HWP
  • 364345 03/08/12 01:32 pzhu 1H05A harp scan
  • 364343 03/08/12 01:29 marcoapc Scalers for run #21962 (right) and #2865 (left)
  • 364338 03/08/12 00:40 M. Cummings Continue with harp scans
  • 364337 03/08/12 00:08 Ole Hansen Swing Shift Summary
  • Wednesday

  • 364336 03/07/12 23:12 Ole Hansen More harp scans underway
  • 364334 03/07/12 22:18 J. Shields [Target] Ramp to 2.5 T
  • 364333 03/07/12 22:11 Ole Hansen Controlled access ends
  • 364332 03/07/12 22:08 J. Shields [Target] 4k alarm set off again, self-corrected
  • 364330 03/07/12 21:05 d. keller [target] flows
  • 364329 03/07/12 21:05 A. Camsonne Harp IH05 withou beam
  • 364328 03/07/12 20:59 jshields [Target] Ramp magnet to zero at request of MCC for short controlled access
  • 364327 03/07/12 20:56 Ole Hansen Short controlled access, target field off
  • 364324 03/07/12 20:44 Ole Hansen More harp scans
  • 364321 03/07/12 19:58 pzhu MCC checking local beam dump right now
  • 364317 03/07/12 19:33 pzhu spot++ in 100nA CW
  • 364313 03/07/12 19:27 pzhu cannot connect hareboot32,happex crate,dvcstlab1 portserver,ts0 in right arm
  • 364312 03/07/12 19:18 pzhu turn off fast raster
  • 364311 03/07/12 19:12 pzhu beam position during last harp scan
  • 364310 03/07/12 19:11 pzhu 1H04 harp scan
  • 364309 03/07/12 19:03 pzhu 1H05A harp scan
  • 364307 03/07/12 18:54 pzhu 1H05A harp scan
  • 364306 03/07/12 18:40 pzhu 1H05A harp scan
  • 364305 03/07/12 18:37 pzhu 1H04 harp scan
  • 364304 03/07/12 18:26 Ole Hansen Target alarm harmless
  • 364303 03/07/12 18:22 Ole Hansen Target alarm
  • 364302 03/07/12 18:17 jpchen Hall C breed-through to Hall A 50 nA
  • 364301 03/07/12 18:09 Ole Hansen COmmencing harp scans
  • 364300 03/07/12 18:02 Ole Hansen MCC is ready for harp scans
  • 364299 03/07/12 17:09 Zhihong Ye Beam Permit
  • 364298 03/07/12 16:45 jmaxwell [Target] Magnet at 2.5T
  • 364297 03/07/12 16:29 Ole Hansen Septa to zero current
  • 364296 03/07/12 16:24 jmaxwell [Target] Ramping Magnet to 2.5T
  • 364295 03/07/12 16:19 jpchen Day shift summary
  • 364292 03/07/12 15:47 jpchen FZ2 current off more than 20 A (5%)
  • 364291 03/07/12 15:42 M. Cummings/ T2 Rate [late entry]
  • 364290 03/07/12 15:01 kamyers [Target] return of the 4K temperature excursion
  • 364289 03/07/12 14:55 Jixie Zhang Beam trajectory for 2.5T, 90 deg target field
  • 364288 03/07/12 12:33 jmaxwell [Target] Magnet at Zero, Leads Locked and Tagged
  • 364287 03/07/12 12:19 kamyers [Target] magnet ramped to zero
  • 364286 03/07/12 12:02 jpchen ion chamber and BPM strip charts during Yves beam test this morning
  • 364285 03/07/12 11:50 jpchen Hall A going to controlled access after a full survey
  • 364284 03/07/12 11:33 kamyers [Target] 4K temp excursions pretty regular
  • 364283 03/07/12 11:25 jpchen beam tuning through chicane with target field changing
  • 364282 03/07/12 11:05 Kalyan Temperature rise in local dump (last night)
  • 364281 03/07/12 10:31 jmaxwell [Target] Second 4K supply Alarm
  • 364280 03/07/12 10:24 Zhihong Ye G2p Status Update
  • 364279 03/07/12 10:19 jpchen Hall back to beam permit for harp scan
  • 364278 03/07/12 10:09 kamyers [Target] 4K supply temperature is climbed to 19K, back to nominal
  • 364277 03/07/12 09:53 jmaxwell [Target] Magnet Ramped to 2.5T
  • 364276 03/07/12 08:38 Kalyan Yves ELOG about beam transport problems
  • 364275 03/07/12 07:22 Zhihong Ye Remain in Control Access before I double check with Yves
  • 364274 03/07/12 03:02 Zhihong Ye G2p Status Update: No beam for rest of the owl shift and Restricted Access Tomorrow to diagonize beamline problems
  • 364273 03/07/12 02:43 Zhihong Ye No Halog posted about when and why septa magnets was off
  • 364272 03/07/12 02:26 M. Cummings [Target] Target Magnet Ramping down
  • 364271 03/07/12 02:21 pzhu turn off septum magnet
  • 364270 03/07/12 02:17 pzhu septum current during owe shift
  • 364269 03/07/12 01:48 pzhu tunning beam trouble
  • 364268 03/07/12 00:51 M. Cummings [Target] target position moved to home
  • 364267 03/07/12 00:51 M. Cummings [Target] Ramping Magnet to 2.5T
  • 364266 03/07/12 00:40 Zhihong Ye Halog halog halog .....
  • 364265 03/07/12 00:34 pzhu
  • 364264 03/07/12 00:30 pzhu hall A generate tool
  • Tuesday

  • 364263 03/06/12 23:55 Ole Hansen Swing Shift Summary
  • 364262 03/06/12 23:41 Ole Hansen Hall Status
  • 364258 03/06/12 21:28 adaq bpm,bcm receiver setting backup
  • 364257 03/06/12 20:03 D. Keller [Target] Moved
  • 364256 03/06/12 19:56 D. Keller D. Keller
  • 364255 03/06/12 19:54 D. Keller [Target] Moved target
  • 364254 03/06/12 19:53 D. Keller [Target] Moved target
  • 364253 03/06/12 19:53 Ole Hansen Cycle Left Q3
  • 364252 03/06/12 19:51 jmaxwell [Target] Magnet Ramped
  • 364251 03/06/12 19:46 Ole Hansen Hall Status
  • 364250 03/06/12 19:45 mhuang LHRS Q3 current scale down study-II
  • 364249 03/06/12 19:28 jmaxwell [Target] Ramping Magnet to 2.5T
  • 364247 03/06/12 18:01 M. Cummings VDC plots from run 21912 (with beam)
  • 364235 03/06/12 16:10 jmaxwell [Target] Magnet Ramp Up Test Complete
  • 364234 03/06/12 15:57 Techs Checklist
  • 364233 03/06/12 14:02 jmaxwell [Target] Starting Magnet Ramp Test Procedure
  • 364232 03/06/12 12:23 jpchen Hall A radiation survey this morning
  • 364231 03/06/12 11:03 A. Camsonne BPM calibration summary
  • 364230 03/06/12 10:50 Zhihong Ye Current Progress Update
  • 364228 03/06/12 09:55 jpchen harp test plots
  • 364227 03/06/12 09:24 Zhihong Ye G2p/Gep Status
  • 364226 03/06/12 08:58 X. Zheng Hall A Tools Screen snapshot
  • 364225 03/06/12 08:47 mhuang/jixie LHRS Q3 scale down optics study
  • 364224 03/06/12 08:14 Chao Septum current scan during the BPM Calibration
  • 364223 03/06/12 08:09 hoyoung Owl Shift Summary March 06, 2012
  • 364222 03/06/12 07:43 hoyoung Restricted access
  • 364221 03/06/12 07:11 Zhihong Ye Ramp down Septa magnet field and request restricted access
  • 364220 03/06/12 07:11 hoyoung Restricted access requested
  • 364219 03/06/12 07:11 marcoapc Replay - fast raster off, slow raster off, no target (pure LHe), delta=0 - run #21931
  • 364217 03/06/12 07:05 marcoapc Replay - Left hrs - L Q3=882.3A - run #2854
  • 364215 03/06/12 07:02 marcoapc Scalers - left L Q3=882.3A - run #2854
  • 364213 03/06/12 07:01 Jixie set LQ3 to 882.3A
  • 364212 03/06/12 07:00 marcoapc Replay - optics - Left hrs - fast raster off, slow raster off, no target (pure LHe), delta=0
  • 364210 03/06/12 06:57 marcoapc Replay - optics - right hrs - fast raster off, slow raster off, no target (pure LHe), delta=0
  • 364207 03/06/12 06:54 marcoapc Scaler - faster raster off, slow raster off, no target (pure Lhe), delta=0
  • 364202 03/06/12 06:50 Jixie Start taking Helium data for optics, delta=0%
  • 364200 03/06/12 06:46 hoyoung Fast raster OFF
  • 364196 03/06/12 06:45 pzhu fast raster calibration
  • 364176 03/06/12 06:31 hoyoung Hall A general tools
  • 364152 03/06/12 06:15 marcoapc Scalers - fast raster calibration - runs #2840 (left) and #21918 (right)
  • 364151 03/06/12 06:14 hoyoung Fast raster settings
  • 364142 03/06/12 06:09 hoyoung Fast raster size of 0.8 x 1.0 mm
  • 364139 03/06/12 06:04 hoyoung 100 nA CW to do fast raster calibration
  • 364138 03/06/12 05:56 M. Yurov [target] Target does not move
  • 364137 03/06/12 05:54 Jixie Set left HRS delta=0%, P=2.251, SeptumI=725.4A
  • 364136 03/06/12 05:42 adaq slow raster calibration
  • 364135 03/06/12 05:35 Jixie Cycle Left Q2,Q3
  • 364134 03/06/12 05:31 hoyoung Slow raster OFF
  • 364125 03/06/12 05:24 marcoapc Scalers - septum Ileft=718.1A and Iright=725.4A - run #21913 (right) and #2835 (left)
  • 364111 03/06/12 05:10 hoyoung Beam position back to 0/0
  • 364109 03/06/12 05:08 pzhu bpm calibration for new run
  • 364105 03/06/12 05:04 marcoapc Replay - Right hrs, septum Iright=725.4A - run #21909
  • 364104 03/06/12 05:03 pzhu beam position fluctuate
  • 364103 03/06/12 05:02 hoyoung 100 nA CW
  • 364102 03/06/12 05:00 marcoapc Replay - Left hrs - septum Ileft=718.1A, 100nA - run #2831
  • 364100 03/06/12 04:59 marcoapc Scalers - septum Ileft=718.1A and Iright=725.4 - runs #2831 (left) and #21909 (right)
  • 364096 03/06/12 04:50 Jixie change septa: Ileft=718.1A, Iright=725.4A
  • 364095 03/06/12 04:48 marcoapc Replay - Left hrs, septum I=703.6A, 8uA - run #2830
  • 364094 03/06/12 04:46 marcoapc Replay - Right hrs, septum I=703.6A, 8uA - run #21908
  • 364093 03/06/12 04:43 marcoapc Scalers - septa I=703.6 - runs #21908 (right) and #2830 (left)
  • 364088 03/06/12 04:22 Jixie change both septa to -3%, 703.6A
  • 364087 03/06/12 04:21 marcoapc Replay - Left hrs, septum=653A - run #21907
  • 364086 03/06/12 04:17 Jixie Change septum current back to 718.1(left) and 725.4(right)
  • 364085 03/06/12 04:16 marcoapc Replay - Left hrs, septum I=653A - run #2829
  • 364084 03/06/12 04:15 hoyoung 8 uA pulse in -4/+4
  • 364083 03/06/12 04:15 marcoapc Scalers - septa in I=653A - runs #2829 (left) and #21907 (right)
  • 364076 03/06/12 04:08 Jixie Change both septa to 653A
  • 364067 03/06/12 04:01 marcoapc Replay - Right hrs, septum I=688A - run #21903
  • 364066 03/06/12 03:59 hoyoung 100 nA CW
  • 364064 03/06/12 03:58 marcoapc Scalers - Septum I=688A - runs #21903 (Right) and #2826 (Left)
  • 364061 03/06/12 03:55 Jixie Change septa currents to 688A
  • 364060 03/06/12 03:52 marcoapc Septum was I=725.4A in run #2825
  • 364059 03/06/12 03:50 marcoapc Replay - Right hrs - run #21902
  • 364058 03/06/12 03:48 adaq bpm setting for new calibration
  • 364057 03/06/12 03:47 hoyoung 8 uA pulse in -4/-4
  • 364056 03/06/12 03:47 marcoapc Replay - Left Hrs - run #2825
  • 364053 03/06/12 03:43 marcoapc Scalers for runs #21902 (Right) and #2825 (Left)
  • 364050 03/06/12 03:40 hoyoung 100 nA CW
  • 364049 03/06/12 03:40 Jixie Change both septa currents to 725.4A
  • 364040 03/06/12 03:31 marcoapc Replay - Left hrs - run #2822
  • 364039 03/06/12 03:29 hoyoung 8 uA pulse in +4/-4
  • 364038 03/06/12 03:28 marcoapc Replay - Right hrs - run #21899
  • 364033 03/06/12 03:23 Jixie Change both septa to 762 (5%)
  • 364030 03/06/12 03:21 marcoapc Scaler - run #21898 (right) and #2821 (left)
  • 364027 03/06/12 03:13 hoyoung 8 uA pulse
  • 364026 03/06/12 03:12 marcoapc Replay - BPM calib - Right, run 21896
  • 364023 03/06/12 03:08 marcoapc Replay - BMP calib - Left, run 2819
  • 364022 03/06/12 03:06 marcoapc Scalers for bpm calib +4,+4
  • 364019 03/06/12 03:02 hoyoung 100 nA CW
  • 364018 03/06/12 03:00 hoyoung BPMA 0/0
  • 364017 03/06/12 02:57 hoyoung 8 uA beam in +4/+4
  • 364010 03/06/12 02:48 hoyoung 100 nA CW
  • 364005 03/06/12 02:32 Pengjia Moved sieve in
  • 364004 03/06/12 02:31 Pengjia disconnect BCM multiplier
  • 364003 03/06/12 02:30 Jixie Change septa currents to 800A
  • 364002 03/06/12 02:27 hoyoung Beam permit
  • 363985 03/06/12 01:14 hoyoung Controlled access
  • 363976 03/06/12 01:04 pzhu bpm calibration done for run before now
  • 363963 03/06/12 00:41 adaq harp 1H04 is getting worse, don't know why
  • Monday

  • 363925 03/05/12 23:01 D. Keller shift summary [harp]
  • 363898 03/05/12 21:31 Kalyan ThirdArm E-plane signals fixed
  • 363887 03/05/12 21:21 Kalyan Third arm raw rates - scaler display
  • 363880 03/05/12 21:13 Kalyan restarted hight voltage GUI
  • 363863 03/05/12 20:36 mhuang Preliminary optics angular calibration
  • 363840 03/05/12 19:35 mhuang/jixie Dp spectrum for carbon and CH2 runs (central hole only)
  • 363833 03/05/12 18:38 pzhu bpm receiver setting backup
  • 363826 03/05/12 18:23 A. Camsonne HARP signal gains
  • 363825 03/05/12 18:05 A. Camsonne Requested 8 uA tuned beam for HARP scan
  • 363824 03/05/12 17:31 rakitha Hall C Results for Hall A Large Charge Asymmetry
  • 363823 03/05/12 17:24 A. Camsonne HARP test conclusions
  • 363822 03/05/12 16:14 mhuang Central hole optics data dp vs VertexZ
  • 363821 03/05/12 15:52 Zhihong Ye Hall-A Tool Screen Snapshot
  • 363820 03/05/12 15:51 Zhihong Ye Septa Magnet Power Supply GUI
  • 363818 03/05/12 15:00 X. Zheng Hall sweep in progress
  • 363817 03/05/12 14:59 jpchen Moved sieve slits out
  • 363816 03/05/12 14:43 Chao/Kalyan Switch the dE1L and E1L signal cable
  • 363813 03/05/12 14:13 Zhihong Ye Hall works mostly done. Will close the hall and ask for beam permit soon.
  • 363812 03/05/12 14:03 folts checklist /tech
  • 363811 03/05/12 13:24 M. Cummings RHRS VDC plots for cosmic run 21852
  • 363810 03/05/12 13:05 A. Camsonne 1H04 HARP signal still noisy
  • 363809 03/05/12 13:00 Zhihong Ye BTA working and Shift Leader need to fill out BTA before the end of shift
  • 363806 03/05/12 12:52 Jack Right Arm VDC high voltage
  • 363805 03/05/12 12:52 A. Camsonne Test signal HARP
  • 363804 03/05/12 12:46 R. Zielinski RHRS s2m double-peak
  • 363803 03/05/12 12:40 R. Zielinski RHRS T2 MLU
  • 363797 03/05/12 11:59 Glamazdin Rack 1H75B15 partialy unpowered
  • 363795 03/05/12 11:42 Zhihong Ye Current Status: No an estimation about when we close the hall
  • 363793 03/05/12 11:17 musson bpm receiver
  • 363783 03/05/12 10:45 A. Camsonne HARP status
  • 363781 03/05/12 10:40 M. Cummings RHRS VDC efficiency from cosmic run 21848
  • 363780 03/05/12 10:32 A. Camsonne Switch HV output to 3 for electron detector HV distribution
  • 363778 03/05/12 10:13 pzhu connect back bpm to happex
  • 363775 03/05/12 09:21 Zhihong Ye Status: Experts are working on Harp in the hall
  • 363774 03/05/12 08:34 A. Camsonne Compton electron detector runs
  • 363773 03/05/12 08:30 Zhihong Ye Chicane Change postponed to Tuesday at 7:30am
  • 363772 03/05/12 08:04 hoyoung Restricted Access
  • 363771 03/05/12 07:59 hoyoung Owl Shift Summary March 04, 2012
  • 363770 03/05/12 07:47 Chao ADC & TDC plots for 3rd arm run 40283
  • 363769 03/05/12 07:24 hoyoung Restricted access requested
  • 363764 03/05/12 07:08 marcoapc VDC efficiency - Rhrs - run#21845
  • 363763 03/05/12 07:01 Zhihong Ye Ramp down Septa Magnet field
  • 363756 03/05/12 06:39 hoyoung 1 uA CW beam for compton
  • 363754 03/05/12 04:55 hoyoung Hall A general tools
  • 363752 03/05/12 00:27 M Friend Compton collimator and lead filter changed
  • Sunday

  • 363751 03/04/12 23:56 pzhu hall A generate tools
  • 363750 03/04/12 23:52 pzhu swing shift summary
  • 363749 03/04/12 23:52 pzhu swing shift summary
  • 363748 03/04/12 23:48 pzhu Re. no bpm information from 2754 to 2768 for left arm
  • 363747 03/04/12 23:34 pzhu no bpm information from 2754 to 2768 for left arm
  • 363744 03/04/12 23:16 M Friend Compton spectrum
  • 363743 03/04/12 23:12 M Friend Compton spectrum
  • 363742 03/04/12 22:41 pzhu right arm vdc hits change from 15:58 to 22:10
  • 363739 03/04/12 22:37 pzhu right arm vdc hits change from 15:58 to 22:10
  • 363720 03/04/12 21:05 pzhu VDC hits and efficiency for run 21837
  • 363710 03/04/12 20:44 R. Zielinski RHRS VDC Plots: Run 21836
  • 363701 03/04/12 19:31 A. Camsonne HARP still broken, will do Compton and Calorimeter runs tonight
  • 363697 03/04/12 19:20 A. Camsonne HARP still have issues
  • 363690 03/04/12 18:52 igorko RHRS VDC efficiency
  • 363689 03/04/12 18:35 Kalyan/Chao Access work on thirdarm
  • 363688 03/04/12 18:33 bogdanw Right arm VDC gas
  • 363684 03/04/12 17:54 Mohamed Hafe Compton_chicane_vacuum_status
  • 363681 03/04/12 17:48 Zhihong Ye Late Entry: Asked for a Control Acces at 4:20pm to fix Harp and VDC ...
  • 363680 03/04/12 17:36 pzhu hall A generator tools
  • 363679 03/04/12 16:25 Chao 3rd arm ADC & TDC plot for run 40278
  • 363678 03/04/12 16:12 Jixie, Chao straight through optics summary, part 4
  • 363676 03/04/12 16:10 igorko scalers run #2754
  • 363675 03/04/12 16:09 igorko replay 2754
  • 363663 03/04/12 15:48 Jixie, Chao Set left HRS momentum to quasi-elastic peak, P=2.228,
  • 363659 03/04/12 15:41 josemans Replay - LHRS 2753
  • 363658 03/04/12 15:40 M. Cummings RHRS VDC
  • 363657 03/04/12 15:37 M. Cummings RHRS VDC
  • 363655 03/04/12 15:34 josemans Scaler - LHRS 2753
  • 363653 03/04/12 15:34 M. Cummings RHRS VDC
  • 363651 03/04/12 15:32 josemans Replay - RHRS 21829
  • 363649 03/04/12 15:31 josemans Scaler - RHRS 21829
  • 363648 03/04/12 15:30 josemans Scaler - RHRS 21829
  • 363645 03/04/12 15:24 Jixie, Chao Fast raster on , 1x1 mm, 100nA on CH2 target
  • 363644 03/04/12 15:23 R. Zielinski [ Target] Moved Target to CH2
  • 363643 03/04/12 15:21 Chao Cycle Q2,Q3 and set P=2.251 SeptumI=725.4 for both arms
  • 363638 03/04/12 14:39 Chao, Jixie, straight through optics run summary, part 3
  • 363637 03/04/12 14:38 josemans Scaler - RHRS 21827
  • 363633 03/04/12 14:32 josemans Scaler - RHRS 21827
  • 363630 03/04/12 14:22 pzhu MCC updated bpm calibration constant
  • 363629 03/04/12 14:21 josemans Scaler - RHRS 21826
  • 363627 03/04/12 14:20 Chao set right delta to 4%
  • 363625 03/04/12 14:18 josemans Replay - LHRS 2747
  • 363623 03/04/12 14:11 Chao Threshold for this run is dE=-420mV E=-300mV
  • 363620 03/04/12 14:09 josemans Scaler - RHRS 21825
  • 363619 03/04/12 14:07 josemans Scaler - LHRS 2747
  • 363618 03/04/12 14:05 Chao set left Delta to 4.5%
  • 363606 03/04/12 13:49 Chao, Jixie changed right delta to 3.5%, P=2.172
  • 363603 03/04/12 13:46 Chao Run 21822 is junk
  • 363601 03/04/12 13:43 Chao, Jixie Set left delat to 4%, P=2.161
  • 363592 03/04/12 13:24 Chao set left delta to 3%
  • 363591 03/04/12 13:22 josemans Scaler - LHRS 2741
  • 363586 03/04/12 13:15 josemans Replay - RHRS 21819
  • 363581 03/04/12 13:08 josemans Replay - LHRS 2738
  • 363580 03/04/12 13:04 josemans Scaler RHRS - 21819
  • 363576 03/04/12 12:58 pzhu bpm receiver setting
  • 363575 03/04/12 12:57 josemans Scaler - LHRS 2738
  • 363573 03/04/12 12:54 josemans Replay - LHRS 2735
  • 363571 03/04/12 12:49 mhuang Q3 setting Right
  • 363564 03/04/12 12:40 Jixie set left delta to 1%,right delta to 2%
  • 363561 03/04/12 12:33 josemans Scaler - LRHS 2735
  • 363559 03/04/12 12:32 Chao Right HRS Sieve Pattern with vertex cut
  • 363555 03/04/12 12:23 josemans Replay - RHRS 21816
  • 363554 03/04/12 12:20 Jixie. Min, straight through optics run summary, part 2
  • 363553 03/04/12 12:19 josemans Scaler - RHRS 21816
  • 363548 03/04/12 12:16 josemans Replay - LHRS 2733
  • 363543 03/04/12 12:06 josemans Scaler - LHRS 2732
  • 363540 03/04/12 12:02 Jixie Turn fast raster on for acceptance calibration
  • 363539 03/04/12 12:02 Jixie change left septum back to 725.4A.
  • 363536 03/04/12 12:00 lerose Q3 setting Right
  • 363533 03/04/12 11:54 josemans replay - LRHS 2730
  • 363531 03/04/12 11:52 Jixie, Min, set left septum to 730.3A
  • 363529 03/04/12 11:50 josemans replay - RHRS 21812
  • 363526 03/04/12 11:45 josemans Scaler - LHRS 2729
  • 363525 03/04/12 11:44 josemans Scaler - LHRS 21812
  • 363521 03/04/12 11:39 josemans replay - RHRS 21811
  • 363519 03/04/12 11:31 josemans Scaler - RHRS 21811
  • 363515 03/04/12 11:27 Jixie, Chao, right delta set to 0%
  • 363512 03/04/12 11:22 Zhihong Ye RQ3 Current Setting ISSUE
  • 363511 03/04/12 11:22 josemans replay - LHRS 2726
  • 363510 03/04/12 11:13 josemans Scaler - LHRS 2726
  • 363507 03/04/12 11:04 Jixie, Min, left delta to 0%
  • 363502 03/04/12 10:40 josemans Scaler - RHRS 21808
  • 363501 03/04/12 10:36 Jixie Zhang, straight through optics run summary, part 1
  • 363498 03/04/12 10:27 josemans Scaler - RHRS 21808
  • 363497 03/04/12 10:26 Jixie, Min, change right delta to -2%, P=2.296, septum I=739.9A
  • 363494 03/04/12 10:24 josemans replay - LHRS 2722
  • 363490 03/04/12 10:06 josemans replay - RHRS 21807
  • 363487 03/04/12 09:56 josemans Scaler - LHRS 2722
  • 363485 03/04/12 09:53 mhuang -4% dp solid angle acceptance
  • 363484 03/04/12 09:53 Jixie change left delta to -1%, P=2.274, septum I=732.7
  • 363480 03/04/12 09:45 josemans replay - LHRS 2720
  • 363477 03/04/12 09:35 R. Zielinski RHRS VDC U1/V1 Problem
  • 363472 03/04/12 09:28 mhuang Central sieve hole rays focal plane distribution
  • 363471 03/04/12 09:22 josemans Scaler - LHRS 2719
  • 363468 03/04/12 09:18 josemans Scaler - LHRS 2718
  • 363467 03/04/12 09:18 josemans Scaler - RHRS 21805
  • 363464 03/04/12 09:10 Jixie change left delta to -2%, right delta to -3.5%
  • 363463 03/04/12 09:05 M. Cummings Hall C flipping HWP
  • 363462 03/04/12 09:01 josemans replay - LHRS 2717
  • 363459 03/04/12 08:58 josemans replay RHRS 21804
  • 363458 03/04/12 08:56 josemans Scaler - LHRS 2717
  • 363457 03/04/12 08:54 josemans Scaler - RHRS 21804
  • 363452 03/04/12 08:34 marcoapc Scalers for Delta Scan=-3%
  • 363450 03/04/12 08:30 Jixie left delta=-3%, P=2.319
  • 363438 03/04/12 08:19 Jixie Right Dipole is ready, delta =-4%
  • 363437 03/04/12 08:14 marcoapc Replay - Optics Lhrs - Delta Scan=-4%
  • 363434 03/04/12 08:04 marcoapc Scalers for Delta Scan=-4%
  • 363431 03/04/12 07:56 marcoapc Scaler for Delta Scan = -4%
  • 363428 03/04/12 07:55 Jixie delta =-4%, P=2.341
  • 363427 03/04/12 07:49 mhuang Septa current tuning results
  • 363426 03/04/12 07:42 marcoapc Replay - Optics Lhrs - Larm momentum 2.246
  • 363425 03/04/12 07:38 marcoapc Replay - Optics Rhrs - Rarm momentum 2.2435
  • 363424 03/04/12 07:37 Jixie Zhang, Cycle Q2,Q3, prepare to do delta scan
  • 363421 03/04/12 07:35 marcoapc Scalers for runs 21800 and 2709 - Rarm momentum 2.2435 and Larm momentum 2.246
  • 363420 03/04/12 07:32 marcoapc Rarm momentum 2.2435 in run 21800
  • 363417 03/04/12 07:29 Jixie Zhang change LHRS momentum to 2.246, RHRS momentum to 2.2435
  • 363416 03/04/12 07:19 marcoapc Optics Rhrs - Q=+1% and LeftDipole=-0.2% RightDipole=-0.3%
  • 363413 03/04/12 07:14 marcoapc Replay - Optics Lhrs - Q=+1% and LeftDipole=-0.2% RightDipole=-0.3%
  • 363412 03/04/12 07:12 marcoapc Scaler for Q=+1% and LeftDipole=-0.2% RightDipole=-0.3%
  • 363409 03/04/12 07:06 jixie zhang change left dipole by -0.2% and right dipole by -0.3%
  • 363408 03/04/12 07:04 marcoapc Replay - Optics Lhrs - Q=+1%
  • 363406 03/04/12 06:57 marcoapc Replay - Optics Rhrs - Q=+1%
  • 363403 03/04/12 06:50 marcoapc Scaler for Q=+1%
  • 363400 03/04/12 06:48 Jixie Zhang changed Q1 to +1%
  • 363398 03/04/12 06:40 marcoapc Scaler to Q1=-1%
  • 363396 03/04/12 06:38 marcoapc CODA rhrs rebooted
  • 363394 03/04/12 06:34 marcoapc Scaler to Q1 -1%
  • 363391 03/04/12 06:26 Jixie Zhang change both Q1 by -1%
  • 363387 03/04/12 06:15 T. Badman focal plane variable plots at septum settings -0.5%,-1%,-2%
  • 363385 03/04/12 06:14 Jixie Zhang set both septa to 725.4A
  • 363380 03/04/12 06:04 Jixie Zhang Set both septa to 730.7A
  • 363376 03/04/12 05:50 mhuang Try different optics matrix
  • 363375 03/04/12 05:48 Jixie Change septa to -0.5%
  • 363371 03/04/12 05:37 marcoapc CODA rhms rebooted
  • 363370 03/04/12 05:34 mhuang Right HRS sieve pattern not so good
  • 363362 03/04/12 05:28 Jixie Zhang Changed septum to -2%
  • 363361 03/04/12 05:26 Jixie Zhang Lower septa current by 1%
  • 363357 03/04/12 05:22 mhuang Sieve pattern!
  • 363352 03/04/12 05:14 marcoapc CODA of the rhrs disconnected on run 21788 - rebooting rac 26
  • 363351 03/04/12 05:09 marcoapc reboot CODA rhrs before 21788
  • 363344 03/04/12 04:52 T. Badman Hall C changing hwp
  • 363341 03/04/12 04:48 Jixie Zhang, Sieve slits are now IN
  • 363340 03/04/12 04:32 jpchen Left septum power off for runs 2682-2698.
  • 363336 03/04/12 04:22 marcoapc reboot left hrs CODA before 2698
  • 363328 03/04/12 04:04 T. Badman [Target] crosshair cup scan
  • 363327 03/04/12 03:59 marcoapc the rhrs scaler t3 and t4 is not ok
  • 363324 03/04/12 03:45 marcoapc change 3rd arm prescale factor
  • 363317 03/04/12 03:37 Jixie Zhang Escorted access to move sieve in 20 minutes
  • 363313 03/04/12 03:31 Jixie Zhang Hall A Tools snapshot
  • 363312 03/04/12 03:29 Jixie zhang Start taking optics data
  • 363310 03/04/12 03:24 pzhu crosshair run
  • 363304 03/04/12 03:11 Zhihong Ye G2p Status Update
  • 363295 03/04/12 02:53 mhuang DB 20110303/db_R/L.vdc.dat update
  • 363260 03/04/12 01:54 Zhihong Ye R-VDC-1 has hardware issues
  • 363255 03/04/12 01:44 Jixie Zhang Cycle Q2,Q3, change HRS momenta to 2.251
  • 363250 03/04/12 01:35 T. Badman [Target] moved target position again
  • 363249 03/04/12 01:28 T. Badman [Target] moved target stick
  • 363248 03/04/12 01:26 Jixie Zhang Change septa current: Ileft=716.76 Iright=711.379
  • 363241 03/04/12 01:13 Jixie Zhang Log every change, please
  • 363234 03/04/12 00:20 M. Cummings pion rejector HV matching
  • 363233 03/04/12 00:14 M. Cummings Hv Settings (New)
  • 363232 03/04/12 00:12 J. Zhang Change half-wave plane
  • 363231 03/04/12 00:07 M. Cummings HV Settings (Old)
  • 363230 03/04/12 00:06 A. Camsonne Can't see beam on HARPs
  • Saturday

  • 363229 03/03/12 23:52 D. Keller target moved
  • 363228 03/03/12 23:44 D. Keller Swing Shift Summary
  • 363225 03/03/12 23:15 pzhu changed bpm bandwidth to 700
  • 363222 03/03/12 23:12 D. Keller target moved to carbon
  • 363221 03/03/12 23:01 pzhu bpm bandwidth changed to 1200
  • 363216 03/03/12 22:57 adaq bpm gain
  • 363202 03/03/12 22:44 pzhu bpm gain
  • 363194 03/03/12 22:31 Chao Large Charge Asymmetry Test Results
  • 363183 03/03/12 22:14 pzhu bpm raw signal 100 nA CW, without raster
  • 363175 03/03/12 22:06 pzhu bpm raw signal with 100nA cw, fast rast on
  • 363171 03/03/12 21:35 A. Camsonne Starting tungsten calorimeter setup for 200 nA run 2665 and 21751
  • 363168 03/03/12 21:26 R. Zielinski Adding Jack to VDC discussion
  • 363153 03/03/12 21:05 Zhihong Ye Hall A Tool Screen Shot
  • 363149 03/03/12 20:30 pzhu asymmetry for run 2659
  • 363148 03/03/12 20:09 Sasha Moller results
  • 363147 03/03/12 20:06 A. Camsonne HARP analyzer fixed
  • 363146 03/03/12 20:06 A. Camsonne HARP analyzer fixed
  • 363145 03/03/12 20:02 A. Camsonne VDC checks
  • 363144 03/03/12 20:01 Kalyan Requested controlled access
  • 363143 03/03/12 19:57 R. Zielinski RHRS U1/V1 Efficiency Low
  • 363142 03/03/12 19:54 R. Zielinski RHRS VDC Hits/Efficiency
  • 363141 03/03/12 19:53 R. Zielinski RHRS VDC Wires/TDC
  • 363140 03/03/12 19:46 Kalyan Back to normal laser settings on HallA IA
  • 363139 03/03/12 19:42 Kalyan Third setting for Hall-A IA
  • 363135 03/03/12 19:40 R. Zielinski LHRS PR2 12-33
  • 363134 03/03/12 19:40 R. Zielinski LHRS PR1 24-33 PR2 0-11
  • 363132 03/03/12 19:39 R. Zielinski LHRS PR1 0-23
  • 363127 03/03/12 19:36 R. Zielinski LHRS VDC Hits/Efficiency
  • 363125 03/03/12 19:35 R. Zielinski LHRS VDC Wires and TDC
  • 363123 03/03/12 19:34 R. Zielinski LHRS Cer Plots
  • 363118 03/03/12 19:33 R. Zielinski LHRS s2m TDC 6-15
  • 363116 03/03/12 19:32 R. Zielinski LHRS s2m ADC 11-15 TDC 0-5
  • 363115 03/03/12 19:31 R. Zielinski LHRS s2m ADC 0-11
  • 363110 03/03/12 19:29 Kalyan First setting for Hall-A IA
  • 363109 03/03/12 19:28 R. Zielinski LHRS S1 Plots
  • 363103 03/03/12 19:18 Kalyan Preparing for larger charge asymmetry run
  • 363098 03/03/12 19:07 R. Zielinski RHRS T2 High-Rate
  • 363096 03/03/12 18:43 Zhihong Ye RHRS S2m Scaler Screen
  • 363095 03/03/12 18:42 Zhihong Ye RHRS S1&Cer Scaler Screen
  • 363094 03/03/12 18:42 Zhihong Ye L-HRS Scaler Scree
  • 363091 03/03/12 18:37 Zhihong Ye Scaler Screen
  • 363083 03/03/12 17:57 R. Zielinski RHRS Cer ADC/TDC
  • 363082 03/03/12 17:56 R. Zielinski RHRS s2m ADC 6-15
  • 363081 03/03/12 17:55 R. Zielinski RHRS s2m TDC 12-15 ADC 0-5
  • 363080 03/03/12 17:49 R. Zielinski RHRS s2m TDC 0-11
  • 363079 03/03/12 17:48 R. Zielinski Forgot a Plot. RHRS PS 12-23L
  • 363078 03/03/12 17:47 R. Zielinski RHRS PreShower 0-23L
  • 363077 03/03/12 17:45 R. Zielinski RHRS PreShower 0-23R
  • 363076 03/03/12 17:44 R. Zielinski RHRS Shower 64-79
  • 363075 03/03/12 17:43 R. Zielinski RHRS Shower 32-63
  • 363074 03/03/12 17:41 R. Zielinski RHRS SH 0-31
  • 363073 03/03/12 17:39 R. Zielinski RHRS S1 Plots
  • 363056 03/03/12 15:57 Kalyan BCM calibration run 2648
  • 363051 03/03/12 15:03 A. Camsonne BCM constants updated with those values
  • 363050 03/03/12 14:58 adaq Previous BCM new receiver constants
  • 363047 03/03/12 14:53 A. Camsonne BCM constants for update
  • 363044 03/03/12 14:16 A. Camsonne Added current EPICS variables to left arm replay
  • 363043 03/03/12 14:04 jpchen Moller Starts
  • 363040 03/03/12 13:59 Jixie Zhang HRS Mometum set to quasi-elastic peak 2.2267
  • 363033 03/03/12 13:50 A. Camsonne Calorimeter run complete - resuming beam 100 nA no raster for BPM checks
  • 363032 03/03/12 13:50 jpchen Tungsten calorimeter calibration done, turned off slow raster
  • 363029 03/03/12 13:43 Jixie Zhang Another executable to calculate septa currents
  • 363028 03/03/12 13:32 mhuang VDC HV tripped from #2597
  • 363026 03/03/12 13:21 A. Camsonne Slow raster on for calorimeter run
  • 363019 03/03/12 13:10 A. Camsonne Quick BCM calibration
  • 363000 03/03/12 12:22 A. Camsonne BCM gain settings
  • 362991 03/03/12 11:38 pzhu bpm tail and double peak
  • 362980 03/03/12 10:26 R. Zielinski [ Target ] Reset Magnet Fill Encoder
  • 362978 03/03/12 10:20 R. Zielinski Turned off VDCs on both arms during BPM test
  • 362971 03/03/12 09:30 R. Zielinski VDC HV Tripped on Both arms
  • 362970 03/03/12 09:27 Zhihong Ye Control Access for Pengjia to fix some BPM connections
  • 362969 03/03/12 09:14 Zhihong Ye G2p Current Status Update+Weekend Plan
  • 362940 03/03/12 07:59 hoyoung Owl Shift Summary March 03, 2012
  • 362939 03/03/12 07:58 hoyoung 200nA Pulse beam with raster off
  • 362938 03/03/12 07:53 pzhu BPM receiver pre gain change
  • 362909 03/03/12 07:31 hoyoung Beam is back
  • 362902 03/03/12 07:02 hoyoung Hall A tools screenshot
  • 362897 03/03/12 06:31 T. Badman [Target] buffer supply valve readout incorrect
  • 362896 03/03/12 06:21 hoyoung Beam Permit
  • 362869 03/03/12 04:20 hoyoung Controlled access
  • 362853 03/03/12 03:51 hoyoung Beam back and BPM 0.0/0.0
  • 362852 03/03/12 03:32 Kalyan/Pengj BPM gains
  • 362851 03/03/12 03:14 hoyoung Waiting for beam for 28 mins
  • 362842 03/03/12 02:43 hoyoung 100 nA CW with raster on
  • 362837 03/03/12 02:15 hoyoung Beam back
  • 362836 03/03/12 01:58 Kalyan double peaks in BPM
  • 362835 03/03/12 01:52 adaq hvgui
  • 362824 03/03/12 01:19 Mohamed Compton_Chicane vacuum status
  • 362821 03/03/12 01:06 Mohamed Hafe Compton cavity power stability
  • 362812 03/03/12 00:46 Mohamed Hafe Compton cavity power restored
  • 362809 03/03/12 00:02 A. Camsonne HARP software has a bug
  • Friday

  • 362795 03/02/12 23:24 T. Badman [Target] target position
  • 362789 03/02/12 23:11 A. Camsonne hac_bcm_average constants updated
  • 362788 03/02/12 23:08 Zhihong Ye Is Target IN?
  • 362787 03/02/12 22:39 pzhu beam tunning
  • 362786 03/02/12 22:34 E. Cisbani Target: He level on nose drops below 15 at 5 uA
  • 362781 03/02/12 22:07 A. Camsonne New BCM receiver rough calibration
  • 362772 03/02/12 21:50 A. Camsonne Beam in Hall A - Ion chamber limit setting up starting
  • 362765 03/02/12 21:29 E. Cisbani Target valve problem fixed
  • 362764 03/02/12 20:48 A. Camsonne Check VDC efficiency regularly
  • 362763 03/02/12 20:44 E. Cisbani Control access to fix a target valve problem
  • 362762 03/02/12 20:40 A. Camsonne Target motion fixed
  • 362761 03/02/12 20:29 V. Sulkosky Left HRS VDC Gas Flow
  • 362754 03/02/12 19:54 musson, cams BPM/BCM Tests and Commissioning for G2P.
  • 362753 03/02/12 19:53 A. Camsonne Target motion non functional
  • 362752 03/02/12 19:52 V. Sulkosky Hall A Main Dump Camera Light Operation
  • 362751 03/02/12 18:44 pzhu first spot_R
  • 362750 03/02/12 18:41 pzhu first spot++
  • 362745 03/02/12 18:00 Chao New g2p library for replay
  • 362744 03/02/12 17:37 E. Cisbani Chamber problem in the beam line
  • 362735 03/02/12 17:16 A. Camsonne Beam in Hall A
  • 362734 03/02/12 15:59 LeRose Spectrometer Setting and Monitoring
  • 362730 03/02/12 13:24 LeRose Quadrupole Cycling (new and improved)
  • 362729 03/02/12 13:15 yawei Hall Sweep Finished
  • 362728 03/02/12 12:55 M. Cummings RHRS plots: shower
  • 362727 03/02/12 12:54 M. Cummings RHRS plots: preshower
  • 362726 03/02/12 12:53 M. Cummings RHRS plots: s2m
  • 362725 03/02/12 12:51 M. Cummings RHRS plots: s1 and cherenkov
  • 362724 03/02/12 12:47 M. Cummings RHRS plots: vdc, again
  • 362723 03/02/12 12:43 M. Cummings RHRS plots: vdc
  • 362722 03/02/12 12:39 musson Hall A BPM(s)
  • 362721 03/02/12 12:29 T. Badman [Target] material in nose
  • 362720 03/02/12 12:22 J. Butler(te Check List
  • 362719 03/02/12 12:21 Kalyan Thirdarm in place
  • 362716 03/02/12 12:01 M. Cummings missing channel 32 on prl2
  • 362713 03/02/12 11:51 M. Cummings LHRS plots: adc rocsync
  • 362712 03/02/12 11:50 M. Cummings LHRS plots: prl1 and prl2
  • 362711 03/02/12 11:48 M. Cummings LHRS plots: s2m
  • 362709 03/02/12 11:47 M. Cummings LHRS plots: s1 and cherenkov
  • 362708 03/02/12 11:44 M. Cummings LHRS Plots: vdc
  • 362707 03/02/12 11:08 R. Michaels cycled Q2,Q3. magnet status
  • 362704 03/02/12 10:40 LeRose Quadrupole Cycling
  • Thursday

  • 362603 03/01/12 18:58 gomez Momentum to Septum currents
  • 362594 03/01/12 17:31 A. Camsonne iochacp2 new network settings
  • 362593 03/01/12 17:29 A. Camsonne iochacp2 previous settings
  • 362588 03/01/12 15:14 brads Missing 'Scope from 2nd floor CH -- please return!
  • 362586 03/01/12 15:14 jie Remote Crate Resets